### v2

# Distilling heterogeneous treatment effects: Stable subgroup estimation in causal inference

Abstract
Recent methodological developments have introduced new black-box approaches to better estimate heterogeneous treatment effects; however, these methods fall short of providing interpretable characterizations of the underlying individuals who may be most at risk or benefit most from receiving the treatment, thereby limiting their practical utility. In this work, we introduce causal distillation trees (CDT) to estimate interpretable subgroups. CDT allows researchers to fit any machine learning model to estimate the individual-level treatment effect, and then leverages a simple, second-stage tree-based model to “distill” the estimated treatment effect into meaningful subgroups. As a result, CDT inherits the improvements in predictive performance from black-box machine learning models while preserving the interpretability of a simple decision tree. We derive theoretical guarantees for the consistency of the estimated subgroups using CDT, and introduce stability-driven diagnostics for researchers to evaluate the quality of the estimated subgroups. We illustrate our proposed method on a randomized controlled trial of antiretroviral treatment for HIV from the AIDS Clinical Trials Group Study 175 and show that CDT out-performs state-of-the-art approaches in constructing stable, clinically relevant subgroups.
11footnotetext: The authors would like to thank Sam Pimentel, Yan Shuo Tan, and participants at the American Causal Inference Conference, Johns Hopkins Biostatistics Causal Inference Seminar, UNC Chapel Hill Biostatistics Seminar for their feedback and comments.22footnotetext: Denotes equal contribution.33footnotetext: Yale University. Email: melody.huang@yale.edu, URL: www.melodyyhuang.com44footnotetext: University of Notre Dame. Email: ttang4@nd.edu, URL: https://tiffanymtang.github.io/55footnotetext: University of California Irvine. Email: anamaria.kenney@uci.edu
Keywords: causal inference, treatment effect heterogeneity, subgroup estimation
1 Introduction
While much of the causal inference literature has focused on estimating an average treatment effect for a specific intervention, researchers are often interested in understanding the underlying treatment effect heterogeneity for a given intervention. For example, in medical settings, researchers are often concerned about potential subsets of units who may be harmed by a particular drug or medical intervention. Similarly, in the social sciences, being able to understand who will benefit or could be potentially harmed by a specific intervention is crucial for evaluating the cost-benefit associated with a specific intervention. Though different approaches have been proposed to consider how to better estimate treatment effect heterogeneity across individuals, being able to characterize subgroups of individuals in a study in an interpretable way remains challenging.
In its simplest form, many subgroup analyses in causal inference focus on pre-specified subgroups based upon prior substantive knowledge. However, it can still be challenging to reason about higher-order interactions between covariates. In the absence of prior knowledge, different tree-based approaches have been proposed to discover new subgroups with heterogeneous treatment effects (e.g., Su et al., 2009; Lipkovich et al., 2011; Loh et al., 2015), of which causal trees (Athey & Imbens, 2016) are arguably the most popular. [^1] Citation count was used as a proxy for popularity. In comparing the citation count of the different papers, as of 2025, the original causal tree paper by Athey & Imbens (2016) has over 2,000 citations, while the alternative methods have anywhere from 100-500 citations. Similar to decision trees (Breiman et al., 1984), causal trees partition the underlying covariate space into groups by finding clusters of individuals who have similar treatment effects. Moreover, unlike linear regression approaches, causal trees output a readily-interpretable partition of subgroups while also accounting for potential high-order interactions.
Unfortunately, a drawback to causal trees is that they can be highly unstable and overfit in the presence of even small amounts of noise (e.g., Cattaneo et al., 2022). Recent literature has proposed alternatives to better model heterogeneous treatment effects with ‘metalearners’—e.g., $T$-learners (e.g., Foster et al., 2011), $S$-learners (e.g., Hill, 2011), $X$-learner (Künzel et al., 2019), $R$-learner (Nie & Wager, 2021), to name a few. These black-box metalearners provide theoretical guarantees in the form of quasi-oracle properties in recovering the individual-level treatment effect; however, they no longer produce a tree-like output and thus cannot be directly used to identify interpretable subgroups.
In the following paper, we propose causal distillation trees (CDT) to stably estimate interpretable subgroups. CDT allows researchers to leverage the power of black-box meta learners, while preserving the interpretability of a simple tree output. CDT is a two-stage learner. The first stage estimates a heterogeneous treatment effect model using a flexible metalearner to generate a predicted individual-level treatment effect for each unit. The second stage then distills the information from the predicted individual-level treatment effect through a decision tree to construct interpretable subgroups. The first step serves as a de-noising step to help mitigate the instabilities that decision trees traditionally suffer from. The second step of ‘distillation’ draws upon existing ideas in machine learning (e.g., Hinton, 2015; Menon et al., 2021; Dao et al., 2021). However, unlike typical machine learning settings, where the primary goal is to improve the second stage learner’s predictive ability, we leverage distillation to improve our ability to estimate interpretable subgroups using the second stage decision tree learner.
Our paper provides several key theoretical and methodological contributions. First, while recent papers have similarly proposed the use of a two-stage learner to estimate subgroups (e.g., Foster et al., 2011; Zhang et al., 2021; Rehill, 2024), CDT is agnostic to what first-stage learner researchers use, allowing for a large degree of flexibility in choosing an informative teacher model. Second, unlike previous work, which has largely relied on simulated evidence, we derive theoretical guarantees, proving that CDT consistently recovers the optimal subgroups. To our knowledge, we are the first paper to prove the theoretical properties in recovering subgroups from distillation. Our theoretical results highlight that distillation can offset the instabilities incurred from using causal trees (or other tree-based approaches) in low signal-to-noise regimes, while preserving the interpretability of these methods.
Finally, we propose a novel model selection procedure to help researchers select a teacher model in CDT. Unlike existing model selection procedures, which compare the goodness-of-fit of different models, our selection procedure is explicitly tailored for estimating subgroups. We introduce a new measure, which we call the Jaccard Subgroup Similarity Index (SSI), that quantifies the similarity of subgroups. Researchers can use the subgroup similarity index to select the teacher model that recovers the most stable subgroups.
The paper is structured as follows. Section 2 presents notation and formalizes the notion of a subgroup in causal inference. In Section 3, we introduce causal distillation trees and derive its theoretical properties. In Section 4, we propose a stability-driven model selection procedure to select a teacher model in CDT. We demonstrate the effectiveness of CDT across extensive simulation settings (Section 5), and an AIDS clinical trial case study (Section 6). Section 7 concludes.
2 Setup and Notation
[…]
A common estimand of interest is the average treatment effect (ATE) for a given study:
$\tau_{\textsc{ATE}}=\frac{1}{n}\sum_{i=1}^{n}\left\{Y_{i}(1)-Y_{i}(0)\right\}.$
In practice, a more policy-relevant quantity is the conditional average treatment effect (CATE), which considers the ATE across subsets of individuals in the study population. Existing literature has recommended researchers leverage their substantive expertise to posit what characteristics could potentially moderate the treatment effect.
The focus of this paper is to consider how to construct interesting subgroups for subsets of individuals. We consider a subgroup as a subset of individuals, defined by a discrete partition in the space of $X$, where the partitions in the space correspond to substantively different treatment effects. In particular, we are interested in characterizing different groups of individuals in a study that have different effects from receiving the same treatment.
We formally define a subgroup $\mathcal{G}^{\prime}_{g}:\mathcal{X}\to\{0,1\}$ as
$\mathcal{G}^{\prime}_{g}(X)=\prod_{j=1}^{p}\mathbbm{1}\left\{X^{(j)}\in R^{(j)}_{g}\right\},$ (1)
where $X^{(j)}$ denotes the $j^{th}$ covariate, and $R^{(j)}_{g}\subseteq\text{supp}(X^{(j)})$. Here, $\mathcal{G}^{\prime}_{g}(X)$ represents a collection of binary decision rules that characterize a subset of individuals in the study (Lipkovich et al., 2017). For example, if $\mathcal{G}^{\prime}_{g}(X)=\mathbbm{1}\{X^{\text{gender}}=\text{Female}\}\cdot\mathbbm{1}\{X^{\text{age}}<35\}$, then $\mathcal{G}^{\prime}_{g}(X)$ corresponds to the subset of female-identifying individuals who are under the age of 35. Throughout the paper, we focus on subgroups that are a discrete partition of the covariate space–i.e., $\sum_{g=1}^{G}\mathcal{G}^{\prime}_{g}(X_{i})=1$ for all units $i=1,\ldots,n$ (i.e., each unit belongs to exactly one subgroup).
Given a subgroup $\mathcal{G}^{\prime}_{g}(X)$, we define $\tau^{(g)}$ as the subgroup average treatment effect:
$\tau^{(g)}:=\mathbb{E}\left[Y(1)-Y(0)\mid\mathcal{G}^{\prime}_{g}(X)=1\right].$ (2)
The primary aim in this paper is to estimate an optimal partition, which maximizes treatment effect heterogeneity across subgroups using the most parsimonious set of rules. Specifically, we aim to construct the most parsimonious partition that minimizes the squared loss between the true individual-level treatment effects and the associated subgroup ATEs:
$\{\mathcal{G}_{g}(X)\}_{g=1}^{G}=\operatorname*{arg\,min}_{\{\mathcal{G}^{\prime}_{g}(X)\}_{g}\in\Omega}\mathbb{E}\left[\left\{\tau_{i}-\sum_{g}\mathcal{G}^{\prime}_{g}(X_{i})\cdot\tau^{(g)}\right\}^{2}\right],$ (3)
where $\Omega$ represents a class of partitions in the covariate space of $X$. Under squared loss, the optimal partition corresponds to a subset in the covariate space of $X$ that maximizes the difference in treatment effect heterogeneity across subgroups. In other words, within a partitioned group, the treatment effect heterogeneity is minimized, thereby maximizing the variance in treatment effects across groups. In practice, researchers can choose alternative loss functions that correspond to what are substantively meaningful partitions in the covariate space. However, the theoretical guarantees we derive utilize the squared loss, and generalizing the results for other loss functions is an open avenue of future research.
Our work is distinct from existing literature in considering subgroup ATEs. For example, alternative quantities in the literature include the sorted group average treatment effect (GATES), which considers the ATE across subsets formed based on the magnitude of the estimated effect (i.e., the ATE across individuals who would most benefit) (e.g., Imai & Li, 2024; Chernozhukov, Demirer, Duflo & Fernandez-Val, 2018; Dwivedi et al., 2020). GATES relies on using a machine learning model to generate predictions of the CATE, and then groups the predicted CATEs by highest to lowest to estimates the ATE across each group. While quantities like GATE provide a way for researchers to evaluate the range of possible treatment effects within a study, it falls short of providing interpretability for the characteristics of the units within each group.
3 Causal distillation trees
3.1 Overview
We propose a method, causal distillation trees (CDT), which stably estimates interpretable subgroups. CDT is a two-stage learner (Figure 1 and Algorithm 1). The first stage learner, referred to as the teacher model, learns an informative model for the individual-level treatment effect using the observed data $\{Y,\;Z,\;X\}$. Using this teacher model, we generate a prediction $\hat{\tau}_{i}^{d}$ for the individual-level treatment effect for all units. In the second stage, we distill the predicted $\hat{\tau}_{i}^{d}$’s into interpretable subgroups by fitting a decision tree, called the student model, to predict the $\hat{\tau}_{i}^{d}$’s from the covariates $X$. The estimated decision rules from the decision tree define a discrete partition, which naturally maps to the collection of subgroups $\left\{\hat{\mathcal{G}}_{1}(X),\;\ldots,\;\hat{\mathcal{G}}_{G}(X)\right\}$ (see Figure A1). Finally, using the estimated partitions, we estimate the subgroup ATE for each subgroup.
Figure 1: Overview of Causal Distillation Trees (CDT). CDT leverages a two-stage procedure, which first fits a teacher model to estimate individual-level treatment effects and then fits a student model (e.g., a decision tree) to distill the estimated individual-level treatment effects and produce interpretable subgroups. The two-stage learner is learned using the training data (blue-gray boxes). Using the estimated subgroups, the subgroup ATEs are honestly estimated with a held-out estimation set (yellow box).
Informally, the first stage learner smooths the individual-level treatment effects by projecting $\tau_{i}$ into the basis space of the covariates $X$. By using the projected version of the individual-level treatment effect instead of $\tau_{i}$, the second stage learner will be able to more stably estimate the subgroups, as the teacher model will have de-noised the outcomes.
Throughout the paper, we use a decision tree—specifically, a classification and regression tree (CART) (Breiman et al., 1984)—as the second-stage learner. While researchers can use alternative models as the second-stage learner (e.g., Bargagli-Stoffi et al., 2020; Wang & Rudin, 2022; Wan et al., 2023), we focus on decision trees because the output maps intuitively to a standard interpretation of a subgroup as a discrete partition of the covariate space. In settings when researchers are interested in a linear CATE, it would be useful to consider alternative second-stage learners.
We also incorporate careful sample splitting in CDT to avoid post-selection bias. Specifically, we randomly split the data into a training set to estimate the subgroups $\left\{\hat{\mathcal{G}}_{1}(X),\ldots,\hat{\mathcal{G}}_{G}(X)\right\}$ and a hold-out estimation set to estimate the subgroup ATEs $\hat{\tau}^{(1)},\ldots,\hat{\tau}^{(G)}$ (see Figure 1). This provides an honest estimate of the subgroup ATE (Athey & Imbens, 2016). In addition, to avoid overfitting to the teacher model, we estimate the $\hat{\tau}_{i}^{d}$’s using out-of-sample procedures (e.g., using out-of-bag samples (Wager & Athey, 2018) or cross-fitting (Chernozhukov, Chetverikov, Demirer, Duflo, Hansen, Newey & Robins, 2018)) in the first stage (details in Algorithm 1).
3.2 Consistency of subgroup estimation: Single rule setting
We now investigate the theoretical properties of CDT. We show that under a set of regularity assumptions, CDT consistently recovers the optimal subgroups, and can improve the rate of convergence in estimating subgroups over standard decision trees. The rate of convergence depends on (1) the smoothness in the outcomes, and (2) how much noise there is in the underlying process. We focus our discussion standard greedy tree algorithms, such as CART. In settings when researchers employ alternative algorithms to construct decision trees, some of the assumptions for consistency can be potentially relaxed.
Throughout, we denote $\mathbb{E}_{n}(A_{i})$ as the expectation of a variable $A_{i}$ over the observed sample (i.e., $\mathbb{E}_{n}(A_{i}):=\frac{1}{n}\sum_{i=1}^{n}A_{i}$), while $\mathbb{E}(A_{i})$ refers to the population-level expectation. Furthermore, for each unit $i$, we define $f_{\tau}(X_{i};k)$ as the prediction generated from a decision tree estimated with $\tau_{i}$, at a single split with threshold $k$ (i.e., $f_{\tau}(X_{i};k):=\mathbbm{1}\{X_{i}\leq k\}\mathbb{E}(\tau_{i}\mid X_{i}\leq k)+\mathbbm{1}\{X_{i}>k\}\mathbb{E}(\tau_{i}\mid X_{i}>k)$). Similarly, we define $f^{d}_{\tau}(X_{i};k)$ as the prediction generated from a decision tree using $\hat{\tau}_{i}^{d}$, at a a single split with threshold $k$ (i.e., $f^{d}_{\tau}(X_{i};k):=\mathbbm{1}\{X_{i}\leq k\}\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\leq k)+\mathbbm{1}\{X_{i}>k\}\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}>k)$).
In this section, we will consider the simplest univariate setting, in which we observe a single covariate and the underlying subgroups are defined by a single rule. Section 3.3 will then generalize the results to show consistency of the estimated subgroups, comprised of multiple rules in the multivariate setting. More formally, let $X\in\mathcal{X}\subseteq\mathbbm{R}$, with the underlying subgroups defined as $\mathcal{G}_{1}(X)=\mathbbm{1}\{X\leq s\}$ and $\mathcal{G}_{2}(X)=\mathbbm{1}\{X>s\}$ for some optimal split threshold $s\in\mathbbm{R}$. We define an optimal split under the squared loss as follows.
Definition 3.1 (Optimal Split).
An optimal split $s\in\mathcal{X}$ satisfies $s:=\operatorname*{arg\,min}_{s^{\prime}\in\mathcal{X}}\mathbb{E}\left[\ell(\tau_{i};X_{i},s^{\prime})\right],$ where $\ell$ is defined as the squared loss function: $\ell(\tau_{i};X_{i},s^{\prime}):=(\tau_{i}-f_{\tau}(X;s^{\prime}))^{2}.$
To ensure that there exists an optimal split, we must assume the following.
Assumption 3.1.
Let $F_{X}$ be the cumulative distribution function of the covariate $X$. Define the points $s_{l}$ and $s_{u}$ such that $10<F_{X}(s_{l})<F_{X}(s_{u})<$ 1. Let $s$ be an optimal split as defined in Definition 3.1. Assume that (i) $F_{X}$ is absolutely continuous with a density $dF_{X}$ that is bounded away from zero and is continuous in a neighborhood of $s$, and (ii) $s$ is unique.
Assumption (3.1) guarantees that there exists subgroups of interest to estimate. Intuitively, the constraint on $F_{X}$ ensures that the squared loss function will attain a minimum value for $s\in[s_{l},s_{u}]$. Furthermore, Assumption (3.1) rules out the possibility that there could be more than one potential split point. In particular, we are implicitly ruling out the setting considered in Cattaneo et al. (2022), in which an optimal split does not exist.
Given that the subgroups exist, we next show that the estimated subgroups $\hat{\mathcal{G}}(X)$ from CDT converge to the true subgroups $\mathcal{G}(X)$. To do so, recall that greedy decision tree algorithms guarantee that within a given sample of size $n$, the estimated split point, denoted $\hat{s}_{n}$, minimizes the within-sample loss (i.e., $\hat{s}_{n}:=\min_{s^{\prime}\in\text{supp}_{n}(X)}\mathbb{E}_{n}[\ell(\tau_{i};X,s^{\prime})]$). Thus, in the hypothetical setting where researchers had access to $\tau_{i}$, the estimated subgroups $\hat{\mathcal{G}}(X)$ from a decision tree would give the optimal partitioning of the covariate space of $X$, within a given sample. To hence establish the consistency of CDT, we must consider two gaps. First, CDT uses the predicted treatment effects $\hat{\tau}_{i}^{d}$ in lieu of the individual-level treatment effect $\tau_{i}$. As such, we must constrain the differences that can arise from distilling the raw treatment effects with the teacher model. Second, we must show that the within-sample optimal partitioning will asymptotically recover the population-level optimal partitioning.
To address the first gap, we restrict our focus to valid teacher models.
Assumption 3.2 (Valid Teacher).
Let $s$ be the population-level optimal split, and define $s^{d}:=\operatorname*{arg\,min}_{s^{\prime}\in\mathcal{X}}\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X_{i},s^{\prime})\right]$. Assume the teacher model $\mathcal{M}$ used to distill the treatment effects into $\hat{\tau}^{d}_{i}$ satisfies $s^{d}=s$. In this case, we say that the teacher model is valid.
Under this validity of the teacher model, the population-level optimal subgroups remain unchanged after distilling the individual-level treatment effects. In other words, the optimal population splits from the predicted individual-level treatment effects $\hat{\tau}_{i}^{d}$ are equivalent to the optimal split with the true individual-level treatment effects $\tau_{i}$. A sufficient, though not necessary, condition for a valid teacher model is a correctly-specified teacher model, such that the residual noise that cannot be explained by the model must be uncorrelated with the covariates $X$ (i.e., exogeneity). In randomized control trials, we expect this to be true if the teacher model is flexible enough and the sample size is sufficiently large. Assumption (3.2) can also be relaxed to consider approximately valid teacher models (see Appendix A.2 for additional discussion).
We further assume that the amount that $f^{d}_{\tau}(X;s^{\prime})$ changes as the split point $s^{\prime}$ changes within a neighborhood of $s$ is bounded.
Assumption 3.3.
Let $\tau_{0}=\frac{1}{2}\left\{\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}<s)+\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\geq s)\right\}$. Define $H(s)=\mathbb{E}\left[(\hat{\tau}^{d}_{i}-\tau_{0})1\{X_{i}\leq s\}\right]$. Let $\mathcal{N}$ be a neighborhood around $s$, where $\mathcal{N}:=(l,u)\subset[s_{l},s_{u}]$. Then, assume the following:
  • $\tau(x)$ is continuous in $(l,s)$ and $(s,u)$, where $\lim_{x\to s^{-}}\tau(x)=\tau(s^{-})\text{ and }\lim_{x\to s^{+}}\tau(x)=\tau(s^{+})$.
  • For $\epsilon>0$, there exists an $21<\alpha<$ 2 where: $\inf_{|s^{\prime}-s|<\epsilon,s\neq s^{\prime}}\frac{|H(s)-H(s^{\prime})|}{|s-s^{\prime}|^{\alpha}}>0$.
  • There exists $0<\eta\leq 1$ such that $\forall~{}\epsilon>0$, $\mathbb{E}\left[(\hat{\tau}^{d}_{i}-\tau_{0})^{2}\mathbbm{1}\{s-\epsilon\leq X_{i}\leq s+\epsilon\}\right]\leq C\epsilon^{2\eta}.$
Assumption (3.3) is a standard assumption in establishing consistency of $M$-estimators (e.g., Van der Vaart, 2000, Theorem 5.52). Informally, $\alpha$ and $\eta$ are parameters that control the smoothness of the underlying function around the optimal split point $s$. The assumption rules out settings where $\hat{\tau}_{i}^{d}$ can take on values of infinity close to the optimal split point $s$. In practice, we expect Assumption (3.3) to generally hold across a large class of data generating processes. For example, when $\hat{\tau}_{i}^{d}$ is continuously differentiable in the neighborhood $\mathcal{N}$ and bounded in $s$, then $\alpha=2$ and $\eta=1/2$. Assumption (3.3) also allows for discontinuities at the split point $s$, so long as $\hat{\tau}_{i}^{d}$ is continuous within the neighborhood around the split point.
Finally, we assume the predicted treatment effects have bounded moments.
Assumption 3.4 (Bounded Moments).
Assume there exists constants $M_{2}<\infty$ and $M_{4}<\infty$ such that $\mathbb{E}\left\{\left(\hat{\tau}_{i}^{d}\right)^{2}\right\}\leq M_{2}$ and $\mathbb{E}\left\{\left(\hat{\tau}_{i}^{d}\right)^{4}\right\}\leq M_{4}$.
With Assumptions (3.1)-(3.4), we can establish convergence.
Proposition 3.1 (Convergence in Subgroup Recovery, for Single Covariate Setting).
Let $\hat{\mathcal{G}}_{1}(X_{i}):=\mathbbm{1}\{X_{i}\leq\hat{s}_{n}\}$ denote the subgroup estimated by CDT. Then, under Assumptions (3.1)-(3.4), CDT will consistently recover the optimal subgroups:
$\mathbb{E}\left[\left\lvert\hat{\mathcal{G}}_{1}(X_{i})-\mathcal{G}_{1}(X_{i})\right\rvert\right]\lesssim\left\lvert\hat{s}_{n}-s\right\rvert=O_{p}(n^{-1/2(\alpha-\eta)}).$
Proposition 3.1 formalizes that the rate of convergence depends directly on how smooth the underlying data generating process is. The rate of convergence in the estimated subgroups inherits a characteristic highlighted in Escanciano (2020), which bridged the gap in convergence rate of the estimated split in a decision tree $\hat{s}_{n}$ between a fully continuous setting (considered in (Banerjee & McKeague, 2007; Buhlmann & Yu, 2002)) and discontinuous settings (considered in (Chan, 1993; Kosorok, 2008)).
Consider the setting in which $\hat{\tau}_{i}^{d}$ is continuously differentiable. This would imply a convergence rate of $O_{p}(n^{-2/3})$. On the other hand, when $\hat{\tau}_{i}^{d}$ is discontinuous at $X=s$, then $\hat{s}-s=O_{p}(n^{-1})$. Intuitively, as the underlying data generating process becomes more discontinuous at the split point $s$, the faster the estimated split $\hat{s}_{n}$ will converge to the true split $s$. While these results have been studied in the decision tree and change point estimation literature (see, for example, Banerjee & McKeague, 2007; Buhlmann & Yu, 2002; Escanciano, 2020), we are, to the best of our knowledge, the first to formally connect the notion of split consistency to subgroup recovery.
Proposition 3.1 provides practical guidance in considering what suitable teacher models should be used to best improve the rate of convergence using distillation. In standard decision tree settings, $\alpha$ and $\eta$ are something inherent to the underlying data generating process. In contrast, under distillation, we are considering the underlying smoothness of not the original outcomes $\tau_{i}$, but the distilled $\hat{\tau}_{i}^{d}$, which are constructed from a teacher model. This implies that when the underlying $\tau(X)$ is smooth around the optimal split point but the teacher model correctly constructs piecewise functions around the partitioning points, then the convergence rate can improve from $O_{p}(n^{-2/3})$ to the much faster rate of $O_{p}(n^{-1})$.
In addition to consistently recovering the subgroups, CDT can further improve the stability of the split estimation by de-noising the individual-level treatment effect and increasing the signal strength in the individual-level treatment effect. To illustrate, in the following example, we consider a setting in which researchers have access to the true individual-level treatment effect $\tau_{i}$ and estimate a decision tree using the covariates $X_{i}$. In practice, this is infeasible, as researchers can never observe $\tau_{i}$, but this comparison serves as a helpful benchmark to compare the potential improvements from distillation.
Example 3.1 (Improving stability of splits with distillation).
Consider a setting where $\tau(X)$ is continuously differentiable and bounded in a neighborhood around $s$. Furthermore, assume we can write the individual-level treatment effect and the predicted treatment effects as a function of $\tau(X_{i})$ and a noise term (i.e., $\tau_{i}=\tau(X_{i})+v_{i}$, and $\hat{\tau}_{i}^{d}=\tau(X_{i})+v_{i}^{d}$), where $\mathbb{E}(v_{i})=\mathbb{E}(v_{i}^{d})=0$, both $\text{var}(v_{i})$ and $\text{var}(v_{i}^{d})$ are finite, and $\text{var}(\tau(X_{i}))>0$. Then, under squared loss and mild regularity assumptions (Assumption (B.3)), the relative asymptotic variance of the estimated splits without distillation $\hat{s}^{orig}_{n}$ and with distillation $\hat{s}_{n}$ is
$\frac{\text{asyvar}(\hat{s}^{orig}_{n})}{\text{asyvar}(\hat{s}_{n})}=\left(\frac{\text{SNR}_{distil}}{\text{SNR}_{original}}\right)^{2+4/3},$
where $\text{SNR}_{distil}=\text{var}(\tau(X))/\text{var}(v^{d})$ and $\text{SNR}_{original}=\text{var}(\tau(X))/\text{var}(v)$.
The results from Example 3.1 follow from noting that the asymptotic distribution of the estimated splits in a linear setting follow Chernoff’s distribution (Buhlmann & Yu, 2002; Groeneboom, 1989). While the variance itself is not straightforward to interpret and depends on an Airy function, the ratio of variances simplifies into a function of the signal-to-noise ratios (derivation in Appendix B). Example 3.1 formalizes the intuition that as the first-stage learner more closely approximates $\tau(X)$ (i.e., the conditional expectation function of $\tau_{i}$, given the pre-treatment covariates), the stability of the estimated splits after distillation also improves. We can hence view the first-stage learner as a de-noising, or smoothing, step that filters the original individual-level treatment effect $\tau_{i}$ to a less noisy representation, in the space of $X$.
3.3 General setting with multiple rules
We next extend the consistency results to the multivariate setting, in which we observe $p$ total covariates (i.e., $X\in\mathbbm{R}^{p}$) and the optimal subgroups, satisfying (3), consist of an arbitrary number of rules with different covariates and different split points. Throughout, for each subgroup $\mathcal{G}_{g}(X)$, we denote the total number of binary decision rules as $r_{g}$.
Under this multivariate setting, estimating consistent subgroups requires not only recovering the correct split points, but also selecting the right covariates to split on. We thus must introduce an additional separability condition, which states that conditional on the previously selected binary decision rules, the differences in loss from constructing a relevant and irrelevant rule must be greater than zero (i.e., separable).
Assumption 3.5 (Separability Condition).
For each subgroup index $g\in\{1,\ldots,G\}$ and rule index $r\in\{1,\ldots,r_{g}\}$, without loss of generality, write the subgroup defined by the $r-1$ previously selected binary decision rules as $\mathcal{G}_{g}^{(r-1)}(X)=\prod_{k=1}^{r-1}\mathbbm{1}\left\{X^{(j_{k})}\lesseqgtr s^{(j_{k})}\right\}$. We assume that there is separability, conditional on the previously selected binary decision rules:
$\min_{a\in A_{r},\;b\in B_{r}}\left\{\mathbb{E}\left[\ell(\tau_{i};X^{(a)},s^{(a)})\mid\mathcal{G}_{g}^{(r-1)}(X)=1\right]-\mathbb{E}\left[\ell(\tau_{i};X^{(b)},s^{(b)})\mid\mathcal{G}_{g}^{(r-1)}(X)=1\right]\right\}=\delta>0,$
where $B_{r}=\{j_{r},\ldots,j_{r_{g}}\}$ denotes the set of remaining relevant subgroup feature indices, $A_{r}=\{1,\ldots,p\}\setminus B_{r}$ is the set of irrelevant subgroup feature indices, $s^{(a)}$, $s^{(b)}$ are the optimal split points conditional on $\mathcal{G}_{g}^{(r-1)}(X)=1$ for $X^{(a)}$, $X^{(b)}$ respectively, and $\mathcal{G}^{(0)}_{g}(X)\equiv 1$.
Informally, Assumption (3.5) states that given the previously constructed rules, whether a rule is relevant or not must be distinguishable through the loss function. The constant $\delta$ represents the minimum difference of the population-level loss associated with splitting on a relevant and irrelevant covariate, conditional on the previous splits. This is intuitively similar to the irrepresentable condition for the Lasso (Zhao & Yu, 2006), which states that irrelevant features cannot be too correlated with relevant ones to achieve model selection consistency. If the losses between relevant and irrelevant rules are tied, then this means the tree will be unable to distinguish between what are relevant and irrelevant rules.
Assumption (3.5) ensures that the underlying subgroups have an optimal substructure so that a greedy second-stage decision tree learner is valid. If Assumption (3.5) is not met, standard greedy tree algorithms like CART can fail to recover the optimal partitions even with infinite amounts of data (see (Tan et al., 2024) for a formal discussion). Recent work has introduced alternative algorithms to construct optimal trees under weaker assumptions, but at a much higher computational cost, and are often restricted to specific settings such as binary features, with limited out-of-sample performance gains relative to greedy trees (e.g., Hu et al., 2019; van der Linden et al., 2024). Given the popularity of greedy search approaches and the computational burden of many optimal tree construction approaches, we focus our discussion on settings in which greedy search algorithms can feasibly construct optimal subgroups.
With separability, we can now extend the results from Proposition 3.1.
Theorem 3.1 (Consistency of subgroup estimation with decision trees).
Let Assumptions (3.1)-(3.3) hold for all relevant covariates, conditional on the previously select decision rules. Additionally, assume bounded moments (Assumption (3.4)) and separability (Assumption (3.5)). Then, for all groups $g\in\{1,\ldots,G\}$:
$\mathbb{E}\left(\left\lvert\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\right\rvert\right)\leq$ $\frac{2r_{g}(C_{\tau}+M\delta)}{\delta}\left\lvert\hat{s}^{(k)}_{n}-s^{(k)}\right\rvert+$
$\qquad\frac{2r_{g}}{\delta}\left\{\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i};X^{(k)},s^{(k)})\right]-\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X^{(k)},s^{(k)})\right]\right\},$
where $k=\operatorname*{arg\,max}_{k^{\prime}}\left\{\left\lvert\hat{s}_{n}^{(k^{\prime})}-s^{(k^{\prime})}\right\rvert+\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i};X^{(k^{\prime})},s^{(k^{\prime})})\right]-\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X^{(k^{\prime})},s^{(k^{\prime})}\right]\right\}$, $M$ is a finite positive constant, and $\delta$ corresponds to the constant in the separability condition. Furthermore, applying Proposition 3.1, $\hat{\mathcal{G}}_{g}(X_{i})\stackrel{{\scriptstyle p}}{{\rightarrow}}\mathcal{G}_{g}(X_{i}).$
The results of Theorem 3.1 follow from first noting that the estimated splits will converge in probability to the optimal split (see Lemma 3.1), and then showing that the within-sample loss will converge in probability to the true population-level loss (see Appendix B) In particular, the rate of convergence of $\hat{\mathcal{G}}_{g}(X_{i})$ to $\mathcal{G}_{g}(X_{i})$ is dependent on (1) the degree of separability in the losses between relevant and irrelevant rules (i.e., $\delta$), and (2) the underlying convergence rate of the sample splits to the true optimal splits (i.e., the rate at which $\left\lvert\hat{s}_{n}^{(k)}-s^{(k)}\right\rvert\to 0$). An immediate implication of Theorem 3.1 is that not only will the tree consistently split at the correct cutoff points, but it will also correctly select the important features that should be used in the splitting criteria.
In practice, the total number of subgroups present is unknown, and can be treated as a tuning parameter. We show empirically that using standard pruning approaches result in consistent subgroup estimation with CDT (Appendix D). Furthermore, if researchers want to constrain the maximum number of rules in a subgroup, Theorem 3.1 guarantees that for a fixed number of rules $r$, CDT gives the optimal partitioning of that cardinality.
To help illustrate the benefits of CDT, we provide a numerical example that compares the performance of CDT with a standard causal tree.
Example 3.2 (Numerical comparison of CDT with causal trees).
Consider a simple data generating process, in which $X_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\text{MVN}(0,I)$ and $\tau_{i}=2\cdot\mathbf{1}\left\{X_{i}^{(1)}>0\right\}-\mathbf{1}\left\{X_{i}^{(2)}<-0.5\right\}+\epsilon_{i}$, where $\epsilon_{i}\sim N(0,\sigma_{\tau}^{2})$ (details in Section 5). We see that the number of estimated subgroups using causal tree increases linearly as a function of the sample size, while CDT correctly recovers the true underlying subgroups. Figure 2 visualizes the results.
Figure 2: Comparing the performance of causal trees versus CDT (using causal forests as the teacher model), measured via (A) number of estimated subgroups as well as (B) number of true positive and (C) number of false positive features used in the estimated subgroups. The oracle number of subgroups, true positives, and false positives are shown as dashed gray lines. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
To summarize, we have shown that under a set of regularity assumptions and a well-specified teacher model, CDT is able to consistently recover the set of optimal subgroups. Furthermore, CDT improves the efficiency of subgroup recovery over existing tree-based approaches by first de-noising the underlying data using the teacher model. Distilling thus allows researchers to exploit the interpretable subgroup structure of decision trees, while offsetting the instabilities that usually adversarially impact tree estimation.
3.4 Estimating subgroup average treatment effects
Next, we consider how to estimate the subgroup ATEs. For simplicity, we focus on an experimental setting, with extensions for estimating the subgroup ATE in observational settings in Appendix A.3. More formally, we assume there is random treatment assignment in the study, such that the treatment indicator $Z$ is independent of $Y(1),Y(0)$, and $X$.
Assumption 3.6 (Random Treatment Assignment).
[…]
We also assume positivity, such that all units have a non-zero probability of receiving treatment.
Assumption 3.7 (Positivity).
There exists some constant $0<\eta\leq 0.5$ such that $\eta<\Pr(Z=1\mid X)\leq 1-\eta.$
Finally, we define a subgroup difference-in-means estimator $\hat{\tau}^{(g)}$ for $g=1,\ldots,G$ as
$\hat{\tau}^{(g)}:=\frac{1}{\sum_{i=1}^{n}Z_{i}\cdot\hat{\mathcal{G}}_{g}(X_{i})}\sum_{i=1}^{n}Z_{i}Y_{i}\hat{\mathcal{G}}_{g}(X_{i})-\frac{1}{\sum_{i=1}^{n}(1-Z_{i})\cdot\hat{\mathcal{G}}_{g}(X_{i})}\sum_{i=1}^{n}(1-Z_{i})Y_{i}\hat{\mathcal{G}}_{g}(X_{i}).$ (4)
We will take a finite-sample perspective, where we consider only the variation that arises from the treatment assignment process. Following existing literature, we condition on a set of valid randomizations—i.e., the set of random treatment allocations, such that there are treatment and control units in each subgroup $g\in\{1,\ldots,G\}$. [^2] Because of the sparsity induced by the causal distillation trees, we expect that the total number of subgroups $G$ should be small, relative to the sample size $n$. As such, we anticipate this restriction will have little impact on the results in practice. See Schochet (2024) and Miratrix et al. (2013) for more discussion. With some abuse of notation, we suppress the explicit conditioning in the results presented next.
Because we are using sample splitting and honest estimation to estimate the subgroups, the estimated subgroups can be treated as a priori defined strata within the study. As such, under Theorem 3.1, the subgroup difference-in-means estimator, defined in Equation (4), provides a consistent estimate of the true subgroup ATEs.
Theorem 3.2 (Consistency and Variance of Subgroup Difference-in-Means Estimator).
Under Assumptions (3.1)-(3.7), the subgroup difference-in-means estimator will be an unbiased estimator for the sample subgroup ATE. Moreover, the variance of $\hat{\tau}^{(g)}$ is
$\text{var}(\hat{\tau}^{(g)})=\frac{1}{n_{g}}\left\{\beta_{g}(1)\text{var}_{g}(Y_{i}(1))+\beta_{g}(0)\text{var}_{g}(Y_{i}(0))-\text{var}_{g}(\tau_{i})\right\},$
where $\beta_{g}(z)=\mathbb{E}\{1/\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})\mathbbm{1}\{Z_{i}=z\}\}$ and $\text{var}_{g}\left\{Y(z)\right\}:=\mathbb{E}\left\{\frac{1}{n_{g}-1}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})(Y_{i}(z)-\right.$ $\left.\overline{Y_{g}(z)})^{2}\right\}$ for $z\in\{0,1\}$.
If researchers are interested in the population-level subgroup ATE, then there will be some finite-sample bias that arises from slippage between the estimated subgroups $\hat{\mathcal{G}}(X)$ and the true optimal subgroups $\mathcal{G}(X)$; however, the subgroup difference-in-means estimator will still be a consistent estimator for the population-level subgroup ATE. (See Appendix A.4 for more discussion.)
From Theorem 3.2, we can also construct a conservative variance estimator using the sample analogs for $\text{var}_{g}(Y_{i}(z))$ for $z\in\{0,1\}$:
$\widehat{\text{var}}(\tau^{(g)})=\frac{1}{n_{g}}\left\{\hat{\beta}_{g}(1)\widehat{\text{var}}_{g}(Y_{i}(1))+\hat{\beta}_{g}(0)\widehat{\text{var}}_{g}(Y_{i}(0))\right\},$
where $\hat{\beta}_{g}(z)=\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})\mathbbm{1}\{Z_{i}=z\}\right)^{-1}$. Miratrix et al. (2013) showed that the gap between $\hat{\beta}_{g}(z)$ and $\beta_{g}(z)$ can be upper bounded as a function of $1/n$. Asymptotically, this will be equivalent to the robust Huber–White variance estimator.
To consider the asymptotic distribution of the subgroup difference-in-means estimator, we must account for the fact that as the sample size increases, the underlying subgroup allocation will also change. We leverage recent work in developing central limit theorem results for finite populations (e.g., Schochet, 2024; Li & Ding, 2017) to show that under mild regularity conditions, the subgroup difference-in-means estimator will be asymptotically normal.
Theorem 3.3 (Asymptotic Normality of Subgroup Difference-in-Means Estimator).
Assume that as the sample size $n\to\infty$, the total number of subgroups $G$ remain fixed. Furthermore, assume the proportion of units in each subgroup converges to a proportion $\pi_{g}^{*}\in(0,1)$ (i.e., $n_{g}/n\to\pi_{g}^{*}$, where $10<\pi_{g}^{*}<$ 1 for all $g\in\{1,\ldots,G\}$ and $\sum_{g=1}^{G}\pi_{g}^{*}=1$), and the proportion of treated units converges to $p_{z}$, where $p_{z}\in(0,1)$ (i.e., $\frac{1}{n}\sum_{i=1}^{n}Z_{i}\to p_{z}$). Then, for $z\in\{0,1\}$ and $g\in\{1,\ldots,G\}$, if the following condition holds:
$\lim_{n\to\infty}\frac{1}{\left(\sum_{i=1}^{n}\mathbbm{1}\{Z_{i}=z\}\right)^{2}}\frac{max_{1\leq i\leq n}\hat{\mathcal{G}}_{g}(X_{i})Y_{i}^{2}(z)}{\text{var}(\hat{\tau}^{(g)})}=0,$ (5)
then the subgroup difference-in-means estimator will converge in distribution to $N(0,1)$:
$\frac{\sqrt{n}\left(\hat{\tau}^{(g)}-\mathbb{E}(\hat{\tau}^{(g)})\right)}{\sqrt{\text{var}(\hat{\tau}^{(g)})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}N(0,1).$
The results of Theorem 3.3 can be viewed as a special case of Schochet (2024), Theorem 1. The condition in (5) is a Lindeberg-type condition, which restricts the tail behavior of the underlying potential outcome distribution and is relatively weak.
Following Theorem 3.3, we can apply standard nonparametric tests of treatment effect heterogeneity (see Appendix A). If researchers are interested in evaluating whether the largest treatment effect across the subgroups is statistically significantly different than the other treatment effects, we caution that a different test is needed due to post-selection bias. See (Wei et al., 2024) for more discussion on valid re-sampling procedures for inference in these settings.
4 Stability-driven teacher model selection
In Section 3, we showed that the benefits of distillation arise from the teacher model’s ability to construct a meaningful basis representation of the individual-level treatment effect using the covariate data. In practice, there are different metalearners that researchers must choose from to estimate the individual-level treatment effect. Traditional model selection procedures cannot be used to evaluate the goodness-of-fit of our teacher models, which predict individual-level treatment effects — quantities that are not observable in practice. Furthermore, checking whether the predicted outcomes match the observed outcomes is insufficient to guarantee that the treatment effects are modeled well (Künzel et al., 2019).
We next introduce a novel model selection procedure, designed specifically for subgroup estimation. Our proposed procedure selects a teacher model based on its ability to stably reconstruct subgroups. Stable reconstruction refers to being able to estimate similar subgroups given various perturbations in the underlying data generating process. We consider more stable teacher models as being more appropriate models for the data, as they demonstrate greater robustness to sampling error (e.g., Yu, 2013; Yu & Kumbier, 2020).
To begin, we must quantify what it means for two subgroups to be ‘similar’ and develop a novel measure, referred to as the Jaccard Subgroup Similarity Index (SSI). Informally, SSI considers two different sets of estimated subgroups $\{\hat{\mathcal{G}}^{(1)},\hat{\mathcal{G}}^{(2)}\}$ and computes the proportion of units that belong to the same subgroup across both partitions. If the units are similarly grouped in the two different partitions, then $\hat{\mathcal{G}}^{(1)}$ and $\hat{\mathcal{G}}^{(2)}$ are similar, and SSI is high. If the units are grouped differently, then $\hat{\mathcal{G}}^{(1)}$ and $\hat{\mathcal{G}}^{(2)}$ are less similar, and SSI will be low.
To formalize, let $\hat{\mathcal{G}}^{(1)}=\{\hat{\mathcal{G}}^{(1)}_{1}(X),\ldots,\hat{\mathcal{G}}^{(1)}_{G}(X)\}$ and $\hat{\mathcal{G}}^{(2)}=\{\hat{\mathcal{G}}^{(2)}_{1}(X),\ldots,\hat{\mathcal{G}}^{(2)}_{G}(X)\}$ denote two different candidate sets of subgroups. Then, for $k\in\{1,2\}$, construct the adjacency matrix $C^{(k)}\in\mathbbm{R}^{n\times n}$ which denotes whether or not two distinct units belong in the same subgroup in $\hat{\mathcal{G}}^{(k)}$. Specifically, for $i,j=1,\ldots,n$, let $C_{ij}^{(k)}=1$ if $X_{i}$ and $X_{j}$ belong to the same subgroup in $\hat{\mathcal{G}}^{(k)}$ and $i\neq j$, and 0 otherwise. We then define the Jaccard Subgroup Similarity Index (SSI) as
$\text{SSI}(\hat{\mathcal{G}}^{(1)},\hat{\mathcal{G}}^{(2)})=\frac{1}{2G}\sum_{\mathcal{H}\in\hat{\mathcal{G}}^{(1)}\cup\hat{\mathcal{G}}^{(2)}}\frac{N_{11}(\mathcal{H})}{N_{01}(\mathcal{H})+N_{10}(\mathcal{H})+N_{11}(\mathcal{H})},$ (6)
where $N_{qr}(\mathcal{H})=\lvert\{(i,j)\in[n]\times[n]:C^{(1)}_{ij}=q,\,C^{(2)}_{ij}=r,X_{i}\in\mathcal{H}(X)\}\rvert$ for each subgroup $\mathcal{H}\in\hat{\mathcal{G}}^{(1)}\cup\hat{\mathcal{G}}^{(2)}$ and $q,r=0,1$. $N_{11}(\mathcal{H})$ is the number of sample pairs that belong to the same subgroup in both partitions ($\hat{\mathcal{G}}^{(1)}$ and $\hat{\mathcal{G}}^{(2)}$), conditioned on the sample pair belonging to subgroup $\mathcal{H}(X)$. $N_{01}(\mathcal{H})+N_{10}(\mathcal{H})$ is the number of sample pairs that belong to subgroup $\mathcal{H}(X)$ in one partition but two different subgroups in the other partition.
The subgroup similarity index is bounded between 0 and 1. When SSI is close to 1 (i.e., when $N_{11}(\mathcal{H})$ is much larger than $N_{01}(\mathcal{H})+N_{10}(\mathcal{H})$), this implies the two sets of subgroups are similar. Figure A2 illustrates the computation of SSI for an example subgroup $\mathcal{H}\in\hat{\mathcal{G}}^{(1)}\cup\hat{\mathcal{G}}^{(2)}$. SSI is similar to the Jaccard index (Jaccard, 1901) used to compare similarities across clusters (Ben-Hur et al., 2001). However, unlike the standard Jaccard index, which weights individual units equally, SSI re-weights units to give equal weight to each subgroup. This penalizes subgroups with very few units, which are typically highly unstable. [^3] We can consider the extreme setting, in which a subgroup contains only a single individual. An estimated subgroup with only one unit would have very little impact on the standard Jaccard index, even though in practice, such a subgroup would not be substantively meaningful and would be undesirable.
Using SSI, we can then assess the similarity of estimated subgroups from each candidate teacher model across repeated data samples (e.g., using bootstrapping). The teacher model that produces the highest SSI across the different bootstrapped iterations corresponds to the most stable teacher model and should be selected (see Algorithm 2 for details).
5 Simulations
In this section, we demonstrate the effectiveness of CDT in accurately estimating subgroups through extensive simulations. We consider three different teacher models: R-learner with boosting (Rboost), causal forest, and Bayesian causal forest (BCF). We benchmark CDT’s performance, relative to popular subgroup estimation approaches — namely, causal trees (Athey & Imbens, 2016), virtual twins (Foster et al., 2011), and interacted linear models, fitted without regularization (i.e., ordinary least squares) and with Lasso ($L_{1}$) regularization (e.g., Imai & Ratkovic, 2013). Implementation and method details can be found in Appendix C. Across the different simulation scenarios, we find that CDT consistently estimates the true subgroup features, thresholds, and CATEs with greater accuracy than existing methods. Moreover, our model selection procedure using the Jaccard SSI provides an effective data-driven approach to choosing an appropriate teacher model for CDT.
Simulation Setup.
We generate a random treatment assignment process, where $Z_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Bernoulli}(1/2)$), and generate $n=500$ samples and $p=10$ covariates from a standard multivariate normal distribution (i.e., $X_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\text{MVN}(0,I)$, where $X_{i}\in\mathbbm{R}^{p}$ and $i=1,\ldots,n$). Furthermore, we simulate three different treatment effect heterogeneity models:
  1. 1.
    ‘AND’ Subgroup DGP: $\tau_{i}=2\mathbf{1}\left\{X_{i}^{(1)}>0\right\}\cdot\mathbf{1}\left\{X_{i}^{(2)}>0.5\right\}+\epsilon_{i}$
  2. 2.
    ‘Additive’ Subgroup DGP: $\tau_{i}=2\mathbf{1}\left\{X_{i}^{(1)}>0\right\}-\mathbf{1}\left\{X_{i}^{(2)}<-0.5\right\}+\epsilon_{i}$
  3. 3.
    ‘OR’ Subgroup DGP: $\tau_{i}=2\mathbf{1}\left\{X_{i}^{(1)}>0\right\}-\mathbf{1}\left\{(X_{i}^{(2)}>0.5)\text{ or }(X_{i}^{(2)}<-0.5)\right\}+\epsilon_{i}$
where $\epsilon_{i}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{\tau}^{2})$, and we define the outcome process to be a function of this underlying treatment effect heterogeneity: $Y_{i}=Z_{i}\cdot\tau_{i}+X_{i}^{(3)}+X_{i}^{(4)}+\nu_{i},$ where $\nu_{i}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,0.1^{2})$. We vary $\sigma_{\tau}$ such that the proportion of variance explained in $\tau(X)$ by the covariates $X$ vary across the values $\{0.2,0.4,0.6,0.8,1\}$. [^4] The proportion of variance explained in $\tau(X)$ by $X$ is defined as $\text{var}(\mathbb{E}[\tau_{i}\mid X])/\text{var}(\tau_{i})\in[0,1]$. To evaluate the effectiveness of each subgroup estimation method, we consider how accurately each method recovers (1) selected subgroup features, (2) estimated subgroup thresholds, and (3) estimated subgroup ATEs (details in Appendix D). We also consider additional data generating processes in Appendix D, and find that the results are largely consistent, even across more complex settings.
Figure 3: Performance of subgroup estimation methods for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via RMSE for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes with linear covariate effects (rows). Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
Subgroup Estimation Results.
Figure 3 visualizes the key simulation results, with more detailed summaries in Appendix D. Across the different simulation settings, we find that the CDT methods are able to accurately estimate the true subgroups in terms of finding the correct features (as measured by $F_{1}$ score in Figure 3A), estimating the correct thresholds to construct the decision rules (Figure 3B), and estimating the true subgroup ATEs (Figure 3C). In contrast, standard subgroup estimation methods (i.e., causal trees and virtual twins) often result in many erroneously constructed subgroups, overfitting to noise features (shown by the high number of false positives). This reinforces our theoretical findings from Section 3 that distillation helps mitigate potential overfitting in the presence of noise. Meanwhile, linear approaches (i.e., linear regression and Lasso) often fail to pick up on existing subgroups (shown by the low number of true positives). The simulations further highlight advantages from distillation in estimating the subgroup ATEs. In addition to generally yielding the lowest RMSE in recovering the subgroup ATEs, we find that CDT is relatively robust to the underlying outcome model’s specification while some existing methods, namely, the linear methods, are highly sensitive to the form of the outcome model (see Figure A3 and Appendix D for details).
Teacher Model Selection Simulations.
To examine our proposed teacher model selection procedure, we computed the Jaccard SSI for Distilled Rboost, Distilled Causal Forest, and Distilled BCF using $B=100$ bootstrap samples and four different choices of tree depths ($d=1,2,3,4$). As a baseline, we also computed the Jaccard SSI for the causal tree.
Figure 4 provides a visual summary of the results under the ‘AND’ subgroup simulation. Under this simulation, we find that regardless of the noise level, Distilled Causal Forest almost uniformly yields the highest subgroup Jaccard stability score, even varying different tree depths. According to our proposed teacher model selection procedure, researchers should thus be using Causal Forest instead of Rboost or BCF as the teacher model. We can directly check the performance of the two teacher models, and find that indeed a higher Jaccard SSI generally corresponds to more accurate subgroup estimation. A similar pattern emerges when examining other simulation scenarios (see Appendix E), thereby demonstrating the efficacy of our stability-driven procedure to select the teacher model.
Figure 4: Under the ‘AND’ subgroup data-generating process with linear covariate effects, we examine (A) the Jaccard SSI across a range of tree depths and (B) the corresponding subgroup estimation accuracy for various subgroup estimation methods (colors) and treatment effect heterogeneity strengths, measured via the proportion of variance explained in $\tau$ by the covariates (columns). Choosing the teacher model in CDT which leads to the highest Jaccard SSI (in this case, the Distilled Causal Forest) generally corresponds to more accurate subgroup estimation.
6 Case Study: AIDS Clinical Trials Group Study 175
To illustrate the stability and interpretability of CDT on real data, we turn to the AIDS Clinical Trials Group Study 175 (ACTG 175). ACTG 175 was a randomized controlled trial to determine the effectiveness of monotherapy compared to combination therapy on HIV-1 infected patients. To enter the study, participants’ CD4 cell counts had to be between 200 to 500 cells per cubic millimeter upon screening. The main outcome of interest was whether a patient reached the primary end point ($Y_{i}=0$), defined by at least a 50% decline in CD4 cell count, the development of acquired immunodeficiency syndrome (AIDS), or death.
Given that combination therapy is already known to slow the progression of HIV-1, our goal is to further characterize subgroups of the population that may experience varying effectiveness of combination therapy using clinical and demographic variables as pre-treatment covariates. Moreover, we specifically focus on the impact of zidovudine and zalcitabine as our treatment compared to a monotherapy regime using zidovudine.
To begin, we perform the proposed model selection procedure to a 50% split of the study, and find that Rboost results in the highest Jaccard SSI across varying tree depths. We then apply CDT, using Rboost as the teacher model, and estimate subgroup ATEs on the remaining 50% (see Table 1 for the estimated treatment effects and descriptions of each subgroup and Figure 5 for a visualization of the resulting tree).
The key covariates in our estimated subgroups are CD8 cell counts, weight at the start of the study, and the amount of previous exposure to anti-retroviral therapy. Importantly, these variables are well-known to influence the progression of HIV-1. CD8 cell counts in particular are used in combination with CD4 counts (McBride & Striker, 2017) to characterize overall immune dysfunction, and a CD4/CD8 ratio above 1 is considered healthy. [^5] For added context on our estimated subgroups, in this study 99% of individuals had CD4 cell counts below 700, following the rule of thumb that a healthy CD4/CD8 ratio is above 1, CD8 values above 700 would place the majority in an unhealthy range. Zidovudine is associated with lipodystrophy (Finkelstein et al., 2015)—a condition that affects how the body stores fat resulting in weight gain in harmful parts of the body (e.g., organs) and loss in others. Thus, a high starting weight can put individuals at higher risk of reaching the primary endpoint. Finally, while previous anti-retroviral therapy can initially help improve CD4 counts, over time, individuals can develop drug resistance and benefit from a different regime.
These three covariates interact across our subgroups, producing varying treatment effects. Those with a higher starting weight benefit more from combination therapy under both lower and higher CD8 counts (i.e., Groups 2 and 3 vs Group 1), though the effect is slightly stronger under a lower (healthier) CD8 count. Notably, it is only when those with a lower starting weight also have the highly risky combination of high CD8 cell counts and substantial previous therapy 666786 days of therapy is in the upper 78th percentile of previous exposure in this study that they benefit significantly from the treatment (Group 4).
To summarize, CDT recovers clinically-relevant subgroups in the AIDS Clinical Trial data that we now know to be substantively important, thus providing a helpful validity check on the usefulness of the proposed method. This illustration also highlights that had researchers utilized CDT when the initial trial was conducted, they could have identified these potential drivers of treatment effect heterogeneity, without knowing them a priori.
We also evaluate the stability of the different methods across bootstraped samples, and find distillation methods both consistently split on the same set of features and that estimated subgroups were overwhelmingly characterized by features with direct clinical relevance. In contrast, the other tree-based approaches split on a variety of different features, depending on the bootstrapped sample, while the linear approaches frequently failed to identify meaningful heterogeneity. See Appendix F.3 for details.
Figure 5: For our case study, we have the resulting decision tree produced by CDT using Rboost as the teacher model. Estimated subgroup ATEs are shown (SE from 1000 bootstrapped samples in parenthesis) and significant effects are outlined in blue.
7 Discussion
We introduced causal distillation trees for interpretable subgroup estimation in causal inference. CDT allows researchers to leverage the flexibility and power of black-box metalearner approaches, while preserving the interpretability of simple decision trees. We prove that CDT allows researchers to consistently estimate the optimal set of subgroups, and we introduce a novel model selection procedure to help researchers use CDT in practice. Furthermore, we have developed an R package causalDT available at https://github.com/tiffanymtang/causalDT, which not only fits CDT but also automatically conducts the subgroup stability and diagnostic checks, thus enabling researchers to interpret their estimated subgroups from CDT alongside important and informative diagnostics in a holistic manner.
There are several promising avenues for future research. First, recent literature in external validity has emphasized the importance of recruiting a sufficiently heterogenous experimental sample that adequately captures the total treatment effect variation across the target population (e.g., Huang & Parikh, 2024; Huang, 2024+). CDT offers an opportunity for researchers to characterize the underlying treatment effect heterogeneity in a given study and to consider how to optimally recruit these subgroups of participants for an externally valid study. Second, CDT provides a very general and flexible two-stage framework. While we focused on CART as the second-stage learner, there are numerous alternative tree-based models (e.g., Loh, 2002; Quinlan, 2014; Hu et al., 2019) and many that have been developed specifically for subgroup estimation (e.g., Loh et al., 2015; Su et al., 2009; Lipkovich et al., 2011; Seibold et al., 2016; Wang et al., 2024), which can be used as second-stage learners in CDT and may further improve performance. Similarly, efforts to further enhance the stability of distillation-based methods (Zhou et al., 2024) can be easily incorporated into the broader CDT framework. Lastly, while this work is focused on the causal inference setting and treatment effect heterogeneity, an extension of CDT could consider prediction models, where there may be variation in the underlying predictive performance of a specific model. In this prediction context, subgroup analyses could help diagnose critical disparities in the performance of predictive models.
Appendix: “Distilling heterogeneous treatment effects: Stable subgroup estimation in causal inference”
Appendix A Extended Discussion
A.1 Pruning
As in the supervised learning setting, decision trees for causal effect estimation (including in the second stage of CDT) are often pruned to avoid overfitting. There are generally two types of pruning: pre-pruning and post-pruning. Pre-pruning refers to techniques or early stopping criteria that are used to stop growing the decision tree during the tree construction process. Common pre-pruning approaches include setting constraints on the maximum depth of the tree, the minimum number of samples per node, or the minimum information gain from making each split. Post-pruning refers to techniques used to remove uninformative nodes from the decision tree after it has already been constructed. For example, in the rpart R package (Therneau et al. 1997), cross-validation is performed to choose the best complexity parameter $\alpha$, where $\alpha$ measures the cost of adding another split to the tree, and all nodes with a cost greater than $\alpha$ are subsequently removed or pruned from the tree.
[…]
A.2 Extended discussion on valid teacher models
[…]
Let $s$ and $s^{d}$ respectively be the optimal splits using the original individual-level treatment effects $\tau_{i}$ and predicted individual-level treatment effects $\hat{\tau}_{i}^{d}$, as defined in Assumption (3.2). Also, define $v_{i}:=\tau_{i}-\hat{\tau}_{i}^{d}$ for each $i=1,\ldots,n$. We can then write
$\tau^{v}(X_{i};s)$ $:=\tau(X_{i};s)-\tau^{d}(X_{i};s)$
$=\mathbbm{1}\{X_{i}\leq s\}\mathbb{E}(v_{i}\mid X_{i}\leq s)+\mathbbm{1}\{X_{i}>s\}\mathbb{E}(v_{i}\mid X_{i}>s)$
$=\mathbb{E}(v_{i}),$
where the last equality follows by exogeneity.
Thus, we have that
$\mathbb{E}\left[\ell(\hat{\tau}_{i}^{d};X_{i},s)\right]$ $=\mathbb{E}\left[\left\{\hat{\tau}_{i}^{d}-\tau^{d}(X_{i};s)\right\}^{2}\right]$
$=\mathbb{E}\left[\left\{\tau_{i}-\tau(X_{i};s)+\tau(X_{i};s)-\tau^{d}(X_{i};s)-v_{i}\right\}^{2}\right]$
$=\mathbb{E}\left[\left\{\tau_{i}-\tau(X_{i};s)-\left(v_{i}-\tau^{v}(X_{i};s)\right)\right\}^{2}\right]$
$=\mathbb{E}\left[\left\{\tau_{i}-\tau(X_{i};s)\right\}^{2}\right]+\mathbb{E}\left[\left\{v_{i}-\tau^{v}(X_{i};s)\right\}^{2}\right]-2\mathbb{E}\left[\left\{\tau_{i}-\tau(X_{i};s)\right\}\cdot\left\{v_{i}-\tau^{v}(X_{i};s)\right\}\right]$
$=\mathbb{E}\left[\left\{\tau_{i}-\tau(X_{i};s)\right\}^{2}\right]+\text{var}(v_{i})-2\underbrace{\mathbb{E}\left[\left\{\tau_{i}-\tau(X_{i};s)\right\}\cdot\{v_{i}-\mathbb{E}(v_{i})\}\right]}_{(*)}.$
Continuing with ()(*)( ∗ ),
$\mathbb{E}\left[\left\{\tau_{i}-\tau(X_{i};s)\right\}\cdot\{v_{i}-\mathbb{E}(v_{i})\}\right]$ $=\mathbb{E}((\hat{\tau}^{d}_{i}+v_{i})(v_{i}-\mathbb{E}(v_{i}))-\tau(X_{i};s)\cdot(v_{i}-\mathbb{E}(v_{i})))$
Applying Law of Iterated Expectations:
$=\mathbb{E}(\mathbb{E}((\hat{\tau}^{d}_{i}+v_{i})(v_{i}-\mathbb{E}(v_{i}))-\tau(X_{i};s)\cdot(v_{i}-\mathbb{E}(v_{i}))\mid X_{i})\mid X_{i})$
Since $\hat{\tau}^{d}_{i}$ and $\tau(X_{i};s)$ are functions of $X_{i}$:
$=\mathbb{E}(\hat{\tau}^{d}_{i}\mathbb{E}(v_{i}-\mathbb{E}(v_{i})\mid X_{i})+\mathbb{E}(v_{i}(v_{i}-\mathbb{E}(v_{i}))\mid X_{i})-$
$\qquad\tau(X_{i};s)\mathbb{E}(v_{i}-\mathbb{E}(v_{i})\mid X_{i}))$
$=\text{var}(v_{i}),$
where the last inequality follows from exogeneity.
Thus,
$\mathbb{E}\left[\ell(\hat{\tau}_{i}^{d};X_{i},s)\right]$ $=\mathbb{E}\left[\ell(\tau_{i};X_{i},s)\right]-\text{var}(v_{i}),$
which implies $s^{d}:=\operatorname*{arg\,min}_{s^{\prime}\in\mathcal{X}}\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X_{i},s^{\prime})\right]=\operatorname*{arg\,min}_{s^{\prime}\in\mathcal{X}}\mathbb{E}\left[\ell(\tau_{i};X,s^{\prime})\right]=:s,$ since $\text{var}(v_{i})$ is not a function of the decision rule split. We have thus shown that exogeneity implies $s=s^{d}$ as desired.
In settings when we do not have a valid teacher model (such that $s^{d}\neq s$), the resulting subgroups will be systematically biased, where
$\mathbb{E}\left[\left\lvert\hat{\mathcal{G}}_{1}(X_{i})-\mathcal{G}_{1}(X_{i})\right\rvert\right]\leq 2M\left\lvert\hat{s}_{n}-s^{d}\right\rvert+\varepsilon,$
where $\varepsilon=2M|s^{d}-s|$.
A.3 Inference for observational settings
While the main manuscript focused on experimental settings, we can similarly estimate the subgroup ATE in observational settings. However, we must account for potential confounding within a given subgroup.
To begin, we must assume conditional ignorability of treatment and outcome.
Assumption A.1 (Conditional ignorability of treatment and outcome).
[…]
This assumption states that given a set of pre-treatment covariates $X$, the treatment assignment process is effectively ‘as-if’ random. While the subgroups themselves are functions of the pre-treatment covariates, they represent a coarsened subset of units in the study. As such, we do not generally expect that all units within a given subgroup have identical covariate values $X$, especially in settings when $X$ is high dimensional. As such, we can no longer rely on just a subgroup difference-in-means estimator, and must adjust for potential confounding of treatment within a subgroup.
Instead of estimating the subgroup difference-in-means, researchers can estimate $\hat{\tau}^{(g)}_{\textsc{w-adj}}$:
$\left\{\hat{\tau}^{(g)}_{\textsc{w-adj}},\hat{\alpha},\hat{\beta}\right\}=\operatorname*{arg\,min}_{\tau,\alpha,\beta}\frac{1}{n_{g}}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\hat{w}_{i}\left\{Y_{i}-\left(\tau Z_{i}+\alpha+\beta^{\top}X_{i}\right)\right\}^{2},$
where $\hat{w}_{i}$ is defined as the inverse propensity score weights:
$\hat{w}_{i}=\begin{cases}1/\hat{e}(X_{i})&\text{if }Z_{i}=1\\ 1/\{1-\hat{e}(X_{i})\}&\text{if }Z_{i}=0\end{cases},$
[…]
Notably, $\hat{\tau}_{\textsc{W-ADJ}}^{(g)}$ can similarly be used in experimental contexts to improve finite-sample performance. In practice, a drawback to subgroup analyses is the relatively small number of observations within each subgroup. In particular, even if the overall study contains a large number of observations, the number of observations available within a subgroup can be relatively limited. As a result, using the standard difference-in-means estimator can result in poor finite-sample performance.
A.4 Extension to infinite population model
In the main manuscript, we focus on the finite-sample setting for considering inference for the subgroup difference-in-means estimator. In settings when researchers are interested in estimating the population-level subgroup ATE, if the sample is a random draw from this larger target population, then we can show that the subgroup difference-in-means estimator will be a consistent estimator. To begin, let $\mathcal{D}$ represent the finite-sample study data. Then:
$\mathbb{E}(\hat{\tau}^{(g)})$ $=\mathbb{E}\left\{\mathbb{E}\left(\hat{\tau}^{(g)}\mid\mathcal{D}\right)\right\}$
$=\mathbb{E}\left[\frac{1}{n_{g}}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})\left\{Y_{i}(1)-Y_{i}(0)\right\}\right]$
$=\mathbb{E}\left[\frac{1}{n_{g}}\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\cdot\left\{Y_{i}(1)-Y_{i}(0)\right\}\right]+\mathbb{E}\left[\frac{1}{n_{g}}\sum_{i=1}^{n}\left\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\right\}\cdot\left\{Y_{i}(1)-Y_{i}(0)\right\}\right]$
Let $K_{\tau}=\max_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left\{Y_{i}(1)-Y_{i}(0)\right\}$:
$\leq\tau^{(g)}+\frac{1}{n_{g}}\sum_{i=1}^{n}K_{\tau}\left\lvert\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\right\rvert$
From Theorem 3.1, $\left\lvert\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\right\rvert\stackrel{{\scriptstyle p}}{{\rightarrow}}0$. As such, $\hat{\tau}^{(g)}\stackrel{{\scriptstyle p}}{{\rightarrow}}\tau^{(g)}$.
The variance of the subgroup difference-in-means estimator would look similar to the variance expression in Theorem 3.3, but with the population-level parameters. Furthermore, when considering the population-level inference, the cross correlation of potential outcomes will disappear. As a result, while the variance estimator will be the same for both inference in a finite-sample setting and an infinite super-population setting, we generally expect more uncertainty associated with inference for the population-level subgroup ATE.
If the study is not a random sample from target population, then we must account for potential distribution shifts between the study sample and the target population. Being able to transport subgroups across different studies to different target populations would require stronger assumptions (see, for example, Zhang et al. 2024).
A.5 Non-parametric test of treatment effect heterogeneity
[…]
$H_{0}:\tau^{(1)}=\ldots=\tau^{(G)}=\tau,$ (7)
for all subgroups $g\in\{1,\ldots,G\}$, with the following estimator:
$\boldsymbol{\hat{\tau}}=(\hat{\tau}^{(1)}-\hat{\tau},\ldots,\hat{\tau}^{(g)}-\hat{\tau})^{\top},$
where $\hat{\tau}:=\frac{1}{\sum_{i=1}^{n}Z_{i}}Y_{i}Z_{i}-\frac{1}{\sum_{i=1}^{n}(1-Z_{i})}Y_{i}(1-Z_{i})$ is the usual difference-in-means estimator across the entire sample. Then, following Theorem 3.3, we can show that under the null hypothesis, the test statistic can be asymptotically approximated using a $\chi^{2}$ distribution. The following corollary formalizes.
Corollary A.1 (Non-Parametric Test of Treatment Effect Heterogeneity).
Under the null hypothesis in (7), with the alternative $H_{\textsc{A}}:\exists~{}\hat{\tau}^{(g)}\neq\tau$, then:
$\boldsymbol{\hat{\tau}}^{\top}\boldsymbol{\Sigma}\boldsymbol{\hat{\tau}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\chi^{2}.$
A.6 Detailed CDT Algorithm
Input: data $\mathcal{D}=\{X,Z,Y\}$, teacher (CATE) model $\mathcal{M}$, student (decision tree) model $m$, training proportion $\pi_{train}$, number of cross-fitting replicates $R$ (if necessary)
1 Randomly split data $\mathcal{D}$ into $\mathcal{D}^{train}=\{X^{train},Z^{train},Y^{train}\}$ and $\mathcal{D}^{est}=\{X^{est},Z^{est},Y^{est}\}$ with probability $\pi_{train}$ and $1-\pi_{train}$, respectively
2 // Stage 1: estimate individual-level treatment effects via teacher model
3 if using model-specific out-of-sample estimation then
4      // e.g., if teacher model has out-of-bag estimation procedure
5      Fit CATE model $\mathcal{M}$ using $\mathcal{D}^{train}$$\rightarrow\widehat{\mathcal{M}}$
6      Use $\widehat{\mathcal{M}}$ to estimate $\tau_{i}$ for each unit $i$ in $\mathcal{D}^{train}$ $\rightarrow\hat{\tau}_{i}^{d}$
7
8      else
9           // do repeated cross-fitting
10           for $r=1,\ldots,R$ do
11                Randomly split $\mathcal{D}^{train}$ into two equally-sized partitions $\mathcal{D}^{train}_{in}$ and $\mathcal{D}^{train}_{out}$
12                Fit CATE model $\mathcal{M}$ using $\mathcal{D}^{train}_{in}$$\rightarrow\widehat{\mathcal{M}}$
13                Use $\widehat{\mathcal{M}}$ to estimate $\hat{\tau}^{d}_{i}$ for each unit $i$ in $\mathcal{D}^{train}_{out}$ $\rightarrow\hat{\tau}^{d}_{i,r}$
14                Reverse roles of $\mathcal{D}^{train}_{in}$ and $\mathcal{D}^{train}_{out}$ and repeat lines 11-12
15
16                end for
17               Average across cross-fits: $\hat{\tau}^{d}_{i}=\sum_{r=1}^{R}\hat{\tau}^{d}_{i,r}$ for each unit $i$ in $\mathcal{D}^{train}$
18
19                end if
20               // Stage 2: estimate subgroups via student model
21                Fit decision tree $m$ on $X^{train}$ to predict $\hat{\tau}_{i}^{d}$’s $\rightarrow$ estimated subgroups $\left\{\hat{\mathcal{G}}_{1}(X),\ldots,\hat{\mathcal{G}}_{G}(X)\right\}$
22                // Honestly estimate subgroup ATEs
23                Estimate $\hat{\tau}^{(g)}$ via (4) for each $g=1,\ldots,G$ using $\mathcal{D}^{est}$
Algorithm 1 Causal Distillation Trees
Figure A1: Given a decision tree (A), its decision splits correspond to binary cuts that partition the covariate space $\mathcal{X}$, shown in (B). Furthermore, the decision splits leading to each terminal node map to the collection of rules that define a subgroup, shown in (C).
A.7 Discussion of Teacher Model Selection Algorithm
We formalize the teacher model selection algorithm in Algorithm 2.
Input: training data $\mathcal{D}^{train}=\{X^{train},Z^{train},Y^{train}\}$, set of candidate teacher (CATE) models $\{\mathcal{M}^{1},\ldots,\mathcal{M}^{L}\}$, student (decision tree) model $m$, number of bootstraps $B$, depth of tree $d$
1 for $l=1,\ldots,L$ do
2      Fit $\mathcal{M}^{l}$ on $\mathcal{D}^{train}$ to obtain estimates of the individual-level treatment effects $\hat{\tau}^{d,l}_{i}$ for each unit in $\mathcal{D}^{train}$ (via lines 2-16 in Algorithm 1)
3      for $b=1,\ldots,B$ do
4           for $t=1,2$ do
5                Get bootstrap sample of $\{(X_{i},\;\hat{\tau}^{d,l}_{i}):i\in\text{training unit}\}$
6                Fit decision tree $m$ on the bootstrapped $X_{i}$’s to predict the bootstrapped $\hat{\tau}^{d,l}_{i}$’s $\rightarrow$ estimated decision tree $\hat{m}^{(b_{t})}$
7                Prune the decision tree $\hat{m}^{(b_{t})}$ to have depth $d$ and obtain estimated subgroups $\hat{\mathcal{G}}^{(b_{t})}$
8
9                end for
10               Compute $J^{l}_{b}:=\mathcal{J}^{subgroup}(\hat{\mathcal{G}}^{(b_{1})},\hat{\mathcal{G}}^{(b_{2})})$
11
12                end for
13
14                end for
15               Select the teacher model $\mathcal{M}^{l^{*}}$ such that $l^{*}=\operatorname*{arg\,max}_{l=1,\ldots,L}\;\;\frac{1}{B}\sum_{b=1}^{B}J^{l}_{b}$
Algorithm 2 Selecting the teacher model for CDT
Notably, the proposed teacher model selection procedure in Algorithm 2 requires choosing the desired tree depth $d$, or equivalently, the number of binary decision rules characterizing each subgroup. This ensures that there are the same number of subgroups (i.e., $2^{d}$) per tree (or partition). Having the same number of subgroups per tree is necessary to ensure that the Jaccard subgroup stability indices are comparable across both bootstraps and teacher models. Otherwise, a teacher model, such as a constant predictor, that results in fewer decision tree splits (or a smaller depth $d$) is inherently going to be more stable than a teacher model that results in more decision tree splits.
Figure A2: Jaccard subgroup similarity index (SSI). (A) Given two candidate partitionings into subgroups, two binary matrices, denoted $C^{(1)}$ and $C^{(2)}$, are constructed, indicating whether or not two distinct units belong in the same subgroup in $\hat{\mathcal{G}}^{(1)}$ and $\hat{\mathcal{G}}^{(2)}$, respectively. (B) $C^{(1)}$ and $C^{(2)}$ are compared to identify the sample pairs which (i) belong to the same subgroup in both partitions (i.e., $N_{11}$ shown as full blue cells), (ii) belong to the same subgroup in one partition and a different subgroup in the other partition (i.e., $N_{01}+N_{10}$ shown as divided blue and gray cells), and (iii) belong to different subgroups in both partitions (i.e., $N_{00}$ shown as full gray cells). (C) Finally, $N_{11}(\mathcal{H})/(N_{01}(\mathcal{H})+N_{10}(\mathcal{H})+N_{11}(\mathcal{H}))$ is computed per subgroup $(\mathcal{H})$ and averaged across all subgroups to obtain the Jaccard SSI.
A.8 Diagnostic Tools for CDT
To help researchers evaluate the quality of the estimated subgroups from CDT, we recommend several important diagnostics.
Prediction accuracy of student model.
First, researchers can assess how well the student model fits the teacher model’s estimated CATEs. We quantify this via the prediction accuracy (e.g., root mean squared error (RMSE)) between the student model predictions and the CATEs, estimated by the teacher model. High concordance between the student model predictions and the teacher model CATEs is a positive sign while low concordance suggests that the student model is not compatible or a good proxy for the teacher model and should lower our degree of trust in the student model’s estimated subgroups.
Distribution of teacher-estimated CATEs.
Secondly, researchers can quantify and visualize the variance or distribution of CATEs, estimated by the teacher model, in each decision tree split in the student model. This is important for understanding both how much heterogeneity occurs within each subgroup (or node) and how much the subgroup CATEs (i.e., the mean CATE per node) change after making each split. A good split (or decision rule) will result in a large change in the mean subgroup CATE and low heterogeneity within the resulting subgroups. In contrast, splits that do not substantively change the mean subgroup CATE nor reduce the amount of heterogeneity within the subgroups can be deemed as uninformative and unnecessary splits.
Number of treated and control units per subgroup.
Thirdly, we recommend reporting the number of treated and control units from the training data that fall in each subgroup. Since the student model only uses $X$ to predict the teacher-estimated CATEs, it is possible that only control or only treated units fall into a student-estimated subgroup. If the training and estimation data splits (i.e., $D^{train}$ and $D^{est}$ from Algorithm 1) are from similar populations, then the estimation data split may similarly yield only control or only treated units (or some highly imbalanced ratio of treated-to-control units) in that subgroup. In these cases, estimation of the subgroup ATE would be highly variable or not possible using (4). It may also suggest that the tree has been grown too deep and that there is currently not sufficient sample size to estimate such a subgroup.
Stability of selected features.
Finally, in addition to evaluating the subgroup Jaccard stability across the different bootstrap samples in Algorithm 2, we can also track the selected features across these bootstraps for each chosen tree depth and plot their distribution (e.g., Figures A10-A15). Similar to the subgroup Jaccard stability, a more stable (or homogeneous) distribution of selected features should elucidate greater trust in that model or tree depth. On the flip side, a very heterogeneous distribution of selected features indicates that the estimated decision tree structure is highly volatile, depending heavily on the training data subset.
Appendix B Proofs and Derivations
B.1 Proof of Proposition 3.1
Proof.
We can re-write the difference in $\hat{\mathcal{G}}_{1}(X_{i})$ and $\mathcal{G}_{1}(X_{i})$ as:
$\mathbb{E}\left[\left\lvert\hat{\mathcal{G}}_{1}(X_{i})-\mathcal{G}_{1}(X_{i})\right\rvert\right]$ $=\mathbb{E}\left[\left\lvert\mathbbm{1}\{X_{i}\leq\hat{s}_{n}\}-\mathbbm{1}\{X_{i}\leq s\}\right\rvert\right]$
$=\mathbb{E}\left[\mathbbm{1}\{X_{i}\leq\hat{s}_{n}\}-\mathbbm{1}\{X_{i}\leq s\}\mid\hat{s}_{n}>s\right]\Pr(\hat{s}_{n}>s)$
$\qquad+\mathbb{E}\left[\mathbbm{1}\{X_{i}\leq s\}-\mathbbm{1}\{X_{i}\leq\hat{s}_{n}\}\mid\hat{s}_{n}\leq s\right]\Pr(\hat{s}_{n}\leq s)$
$=2\left\lvert\mathbb{E}\left[\mathbbm{1}\{X_{i}\leq\hat{s}_{n}\}-\mathbbm{1}\{X_{i}\leq s\}\right]\right\rvert$
$=2\left\lvert F_{X}(\hat{s}_{n})-F_{X}(s)\right\rvert$
Because $F_{X}$ is continuously differentiable by Assumption (3.1), there must exist a positive, real-valued constant $M$ such that $f_{X}(x)\leq M<\infty$:
$\leq 2M\left\lvert\hat{s}_{n}-s\right\rvert.$
[…]
$\mathbb{E}\left[\left\lvert\hat{\mathcal{G}}_{1}(X_{i})-\mathcal{G}_{1}(X_{i})\right\rvert\right]=O_{p}(n^{-1/2(\alpha-\eta)}).$
B.2 Proof of Theorem 3.1
To prove Theorem 3.1, we will use two useful lemmas. Lemma B.1 states that the sample loss function will converge in probability to the population-level loss.
Lemma B.1 (Bounded Loss).
Without loss of generality, assume $s^{(k)}_{n}<s$. Then for $k\in\{1,...,p\}$:
$\left|\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i},X^{(k)};\hat{s}^{(k)}_{n})\right]-\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X^{(k)},s^{(k)})\right]\right|\leq C^{(k)}_{\tau}\times\left|\hat{s}^{(k)}_{n}-s^{(k)}\right|+\bigg{\lvert}\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i};X,s)\right]-\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X,s)\right]\bigg{\rvert}$
where $C^{(k)}_{\tau}=2\max_{c\in[\hat{s}^{(k)}_{n},s^{(k)}]}\mathbb{E}(\hat{\tau}^{d}_{i}\mid X^{(k)}_{i}<c)\times\mathbb{E}_{n}\left|\tau^{d}(X^{(k)}_{i};\hat{s}^{(k)}_{n})-\hat{\tau}^{d}_{i}\right|$.
An immediate implication of Lemma B.1 is that the within-sample loss at the sample-optimal split $\hat{s}_{n}$ will converge in probability to the population-level loss at the population-optimal split point $s$.
Lemma B.2.
Let $X^{(a)}$ be a relevant covariate that has been omitted, whereas $X^{(b)}$ is an irrelevant covariate, erroneously chosen. We can upper bound the probability of the within-sample loss for an irrelevant covariate $X^{(b)}$ being less than a relevant covariate $X^{(a)}$:
$\Pr$ $\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}_{n}^{(b)})\right\}<\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}\right]$
$\leq\frac{2}{\delta_{ba}}\mathbb{E}\left[\left|\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right|\right],$
where $\delta_{ba}:=\mathbb{E}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},s^{(b)})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},s^{(a)})\right\}$.
Intuitively, if the population-level losses between an irrelevant and relevant covariate is relatively small (i.e., $\delta_{ba}$ is small), then it will take longer for the upper bound to approach zero.
We can now prove Theorem 3.1. Let $s^{(k)}$ be the population-level optimal split for covariate $k$ and $\hat{s}^{(k)}_{n}$ be the estimated optimal split for covariate $k$ using a sample size of $n$.
Simple Case: Single rule.
We consider the simple setting in which there is a single rule. Assume $X^{(a)}$ is the covariate that the estimated subgroup chooses to construct a rule with, while $X^{(b)}$ is the covariate that the true subgroup uses to construct a rule. We can bound the difference between the estimated and optimal subgroup:
$\mathbb{E}$ $\left[\left\lvert\hat{\mathcal{G}}_{1}(X)-\mathcal{G}_{1}(X)\right\rvert\right]$
$=\mathbb{E}\left[\left\lvert\mathbbm{1}\{X^{(a)}<\hat{s}_{n}^{(a)}\}-\mathbbm{1}\{X^{(b)}<s^{(b)}\}\right\rvert\right]$
$\leq\underbrace{\mathbb{E}\left[\left\lvert\mathbbm{1}\{X^{(a)}<\hat{s}_{n}^{(a)}\}-\mathbbm{1}\{X^{(b)}<\hat{s}_{n}^{(b)}\}\right\rvert\right]}_{(1)}+\underbrace{\mathbb{E}\left[\left\lvert\mathbbm{1}\{X^{(b)}<\hat{s}_{n}^{(b)}\}-\mathbbm{1}\{X^{(b)}<s^{(b)}\}\right\rvert\right]}_{(2)}$ (8)
First, consider Equation (8)-(1):
$\mathbb{E}$ $\left[\left\lvert\mathbbm{1}\{X^{(a)}<\hat{s}_{n}^{(a)}\}-\mathbbm{1}\{X^{(b)}<\hat{s}_{n}^{(b)}\}\right\rvert\right]$
$=\mathbb{E}\left[\left\lvert\left.\mathbbm{1}\{X^{(a)}<\hat{s}_{n}^{(a)}\}-\mathbbm{1}\{X^{(b)}<\hat{s}_{n}^{(b)}\}~{}\right|~{}X^{(a)}\neq X^{(b)}\right\rvert\right]\Pr(X^{(a)}\neq X^{(b)})$
$=\mathbb{E}\left[\left\lvert\left.\mathbbm{1}\{X^{(a)}<\hat{s}_{n}^{(a)}\}-\mathbbm{1}\{X^{(b)}<\hat{s}_{n}^{(b)}\}~{}\right|~{}X^{(a)}\neq X^{(b)}\right\rvert\right]$
×$\qquad\times\Pr\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}_{n}^{(a)})\right\}<\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}_{n}^{(b)})\right\}\right]$
$\leq\Pr\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}_{n}^{(a)})\right\}<\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}_{n}^{(b)})\right\}\right],$
where the first equality follows from Law of Total Expectation, and the second equality follows from noting that the covariate chosen for the first split will differ from the population-level split covariate if the within-sample loss for covariate $X^{(a)}$ is lower than the within-sample loss for covariate $X^{(b)}$. The final inequality follows from the fact that $\mathbb{E}\left[\left\lvert\left.\mathbbm{1}\{X^{(a)}<\hat{s}_{n}^{(a)}\}-\mathbbm{1}\{X^{(b)}<\hat{s}_{n}^{(b)}\}~{}\right|~{}X^{(a)}\neq X^{(b)}\right\rvert\right]$ must be bounded between 0 and 1. In settings when $X^{(a)}$ and $X^{(b)}$ are very correlated, then it could be the case that in fact, $\mathbb{E}\left[\left\lvert\left.\mathbbm{1}\{X^{(a)}<\hat{s}_{n}^{(a)}\}-\mathbbm{1}\{X^{(b)}<\hat{s}_{n}^{(b)}\}~{}\right|~{}X^{(a)}\neq X^{(b)}\right\rvert\right]$ is quite close to zero as the set of units for whom $\mathbbm{1}\{X^{(a)}<\hat{s}_{n}^{(a)}\}$ and $\mathbbm{1}\{X^{(b)}<\hat{s}_{n}^{(b)}\}$ may be quite similar.
Equation (8)-(2) can be bounded by applying the results from Proposition 3.1:
$\mathbb{E}\left[\left\lvert\mathbbm{1}\{X^{(b)}<\hat{s}_{n}^{(b)}\}-\mathbbm{1}\{X^{(b)}<s^{(b)}\}\right\rvert\right]$ $\leq 2M\left\lvert\hat{s}_{n}^{(b)}-s^{(b)}\right\rvert,$
for some constant $M>0$.
Extending to $r$ total rules:
For $k=1,\ldots,r$, assume $X^{(a_{k})}$ is the covariate that the estimated subgroup chooses to construct a rule with at the $k^{th}$ step, while $X^{(b_{k})}$ is the covariate that the true subgroup uses to construct a rule at the $k^{th}$ step.
$\mathbb{E}$ $\left[\left\lvert\prod_{k=1}^{r}\mathbbm{1}\{X^{(a_{k})}<\hat{s}_{n}^{(a_{k})}\}-\prod_{k=1}^{r}\mathbbm{1}\{X^{(b_{k})}<\hat{s}_{n}^{(b_{k})}\}\right\rvert\right]$
$\leq$ $\Pr(X^{(a_{1})}\neq X^{(b_{1})})+\Pr(X^{(a_{1})}=X^{(b_{1})})\cdot\Pr(X^{(a_{2})}\neq X^{(b_{2})}\mid X^{(a_{1})}=X^{(b_{1})})+\ldots$
$+\Pr(X^{(a_{1})}=X^{(b_{1})},\ldots,\Pr(X^{(a_{r-1})}=X^{(b_{r-1})}))$
×$~{}~{}~{}\times\Pr(X^{(a_{r})}\neq X^{(b_{r})}\mid X^{(a_{1})}=X^{(b_{1})},\ldots,X^{(a_{r-1})}=X^{(b_{r-1})})$
$\leq$ $\sum_{k=1}^{r}\Pr\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a_{k})},\hat{s}^{(a_{k})}_{n})\right\}<\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b_{k})},\hat{s}^{(b_{k})}_{n})\right\}\right]+2M\left\lvert\hat{s}^{(b_{k})}_{n}-s^{(b_{k})}\right\rvert$
Let $k^{*}:=\operatorname*{arg\,max}_{k^{\prime}}\left\lvert\hat{s}_{n}^{(k^{\prime})}-s^{(k^{\prime})}\right\rvert+\mathbb{E}_{n}\left[\ell(\hat{\tau}_{i}^{d};X^{(k^{\prime})},s^{(k^{\prime})})\right]-\mathbb{E}\left[\ell(\hat{\tau}_{i}^{d};X^{(k^{\prime})},s^{(k^{\prime})}\right]$, and $\delta=\min\delta_{ba}$. Applying Lemma B.1 and B.2:
$\leq$ $\frac{2r}{\delta}\mathbb{E}\left\{\left\lvert\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i},X^{(k^{*})};\hat{s}_{n}^{(k^{*})})\right]-\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X^{(k^{*})},s^{{(k^{*})}})\right]\right\rvert\right\}+2Mr\left\lvert\hat{s}^{(k^{*})}_{n}-s^{(k^{*})}\right\rvert$
$\leq$ $\frac{2rC_{\tau}}{\delta}\left\lvert\hat{s}^{(k^{*})}_{n}-s^{(k^{*})}\right\rvert+\frac{2r}{\delta}\left(\mathbb{E}_{n}\left[\ell(\hat{\tau}_{i}^{d};X^{(k^{*})},s^{(k^{*})})\right]-\mathbb{E}\left[\ell(\hat{\tau}_{i}^{d};X^{(k^{*})},s^{(k^{*})})\right]\right)+2Mr\left\lvert\hat{s}^{(k^{*})}_{n}-s^{(k^{*})}\right\rvert$
$=$ $\underbrace{\frac{2r(C_{\tau}+M\delta)}{\delta}\left\lvert\hat{s}^{(k^{*})}_{n}-s^{(k^{*})}\right\rvert}_{(1)}+\underbrace{\frac{2r}{\delta}\big{\{}\mathbb{E}_{n}\left[\ell(\hat{\tau}_{i}^{d};X^{(k^{*})},s^{(k^{*})})\right]-\mathbb{E}\left[\ell(\hat{\tau}_{i}^{d};X^{(k^{*})},s^{(k^{*})})\right]\big{\}}}_{(2)}.$
Term (1) represents the gap between an estimated split and true optimal split. Term (2) is the gap between the empirical loss and the population-level loss. From Proposition 3.1, term (1) converges to zero at rate $O_{p}(n^{-1/2(\alpha-\eta)})$. All that remains is to show that the second term also conevrges to zero.
To begin, we re-write term (2) as:
$\mathbb{E}_{n}$ $\left[\ell(\hat{\tau}^{d}_{i};X,s)\right]-\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X,s)\right]$
$=\mathbb{E}_{n}\left\{\left(\hat{\tau}^{d}_{i}\right)^{2}\right\}-2\mathbb{E}_{n}\left\{\hat{\tau}^{d}_{i}\cdot\hat{f}^{d}_{\tau}(X;s^{\prime})\right\}+\mathbb{E}_{n}\left\{\hat{f}^{d}_{\tau}(X;s^{\prime})^{2}\right\}$
$\quad-\left[\mathbb{E}\left\{\left(\hat{\tau}^{d}_{i}\right)^{2}\right\}-2\mathbb{E}\left\{\hat{\tau}^{d}_{i}\cdot f^{d}_{\tau}(X;s^{\prime})\right\}+\mathbb{E}\left\{f^{d}_{\tau}(X;s^{\prime})^{2}\right\}\right],$
where $\hat{f}^{d}_{\tau}(X;s^{\prime}):=\mathbbm{1}\{X_{i}\leq s\}\mathbb{E}_{n}(\hat{\tau}_{i}^{d}\mid X_{i}\leq s)+\mathbbm{1}\{X_{i}>s\}\mathbb{E}_{n}(\hat{\tau}_{i}^{d}\mid X_{i}>s\}$ is the sample analog to $\tau^{d}(X,s)$. Then:
$=\underbrace{\mathbb{E}_{n}\left\{\left(\hat{\tau}^{d}_{i}\right)^{2}\right\}-\mathbb{E}\left\{\left(\hat{\tau}_{i}^{d}\right)^{2}\right\}}_{(a)}-\underbrace{\left[\mathbb{E}_{n}\left\{\hat{f}^{d}_{\tau}(X;s^{\prime})^{2}\right\}-\mathbb{E}\left\{f^{d}_{\tau}(X;s^{\prime})^{2}\right\}\right]}_{(b)}.$
We will now tackle each term. For term (a), we can directly apply Chebyshev’s Inequality, where for $c>0$:
$\Pr\left(\left\lvert\mathbb{E}_{n}\left\{\left(\hat{\tau}^{d}_{i}\right)^{2}\right\}-\mathbb{E}\left\{\left(\hat{\tau}^{d}_{i}\right)^{2}\right\}\right\rvert\geq\frac{c}{2r}\right)\leq\frac{4M_{4}r^{2}}{nc^{2}},$
where the inequality follows from the assumption of bounded 4th moment. Then, taking $c=2\sqrt{3M_{4}}rt/\sqrt{n}$:
$\Pr\left(\left\lvert\mathbb{E}_{n}\left\{\left(\hat{\tau}^{d}_{i}\right)^{2}\right\}-\mathbb{E}\left\{\left(\hat{\tau}^{d}_{i}\right)^{2}\right\}\right\rvert\geq\frac{\sqrt{3M_{4}}t}{\sqrt{n}}\right)\leq\frac{1}{3t^{2}}.$
For term (b):
$\mathbb{E}_{n}$$\left\{\hat{f}^{d}_{\tau}(X;s^{\prime})^{2}\right\}-\mathbb{E}\left\{f^{d}_{\tau}(X;s^{\prime})^{2}\right\}$
$=$ $\mathbb{E}_{n}\left[\mathbbm{1}\left\{X_{i}\leq s\right\}\right]\cdot\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)^{2}+\mathbb{E}_{n}\left[\mathbbm{1}\left\{X_{i}>s\right\}\right]\cdot\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}>s)^{2}$
$-\left(\mathbb{E}\left[\mathbbm{1}\left\{X_{i}\leq s\right\}\right]\cdot\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)^{2}+\mathbb{E}\left[\mathbbm{1}\left\{X_{i}>s\right\}\right]\cdot\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}>s)^{2}\right)$
$=$ $\left(\mathbb{E}_{n}\left[\mathbbm{1}\left\{X_{i}\leq s\right\}\right]-\mathbb{E}\left[\mathbbm{1}\left\{X_{i}\leq s\right\}\right]\right)\cdot\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)^{2}$
$+\left(\mathbb{E}_{n}\left[\mathbbm{1}\left\{X_{i}>s\right\}\right]-\mathbb{E}\left[\mathbbm{1}\left\{X_{i}>s\right\}\right]\right)\cdot\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}>s)^{2}$
$-\bigg{(}\mathbb{E}\left[\mathbbm{1}\left\{X_{i}\leq s\right\}\right]\cdot\left\{\mathbb{E}(\hat{\tau}_{i}^{d}\mid X_{i}\leq s)^{2}-\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)^{2}\right\}+$
$~{}~{}~{}~{}~{}~{}\mathbb{E}\left[\mathbbm{1}\left\{X_{i}>s\right\}\right]\cdot\left\{\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}>s)^{2}-\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}>s)^{2}\right\}\bigg{)}$
Let $p_{n}(s)=\mathbb{E}_{n}\left[\mathbbm{1}\{X_{i}\leq s\}\right]$, and define $M_{1}=\max_{i}\hat{\tau}^{d}_{i}$. Then:
$\leq$ $2\left\lvert p_{n}(s)-p(s)\right\rvert\cdot M_{1}^{2}+p(s)\cdot\left\lvert\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)^{2}-\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)^{2}\right\rvert$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+(1-p(s))\cdot\left\lvert\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}>s)^{2}-\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}>s)^{2}\right\rvert.$
Then, applying Chebyshev’s inequality again, for all $c>0$:
$\Pr\left(\left|\mathbb{E}_{n}\left[\mathbbm{1}\left\{X_{i}\leq s\right\}\right]-\mathbb{E}\left[\mathbbm{1}\left\{X_{i}\leq s\right\}\right]\right|\geq\frac{c}{4M_{1}^{2}r}\right)$ $\leq\frac{16p_{n}(s)\cdot(1-p_{n}(s))\cdot M_{1}^{2}r}{nc^{2}}\leq\frac{4M_{1}^{2}r}{nc^{2}},$
where $M_{1}<\infty$ is a constant, by the bounded first moment assumption.
Setting $c=\sqrt{12r}tM_{1}/\sqrt{n}$:
$\Pr\left(\left|\mathbb{E}_{n}\left[\mathbbm{1}\left\{X_{i}\leq s\right\}\right]-\mathbb{E}\left[\mathbbm{1}\left\{X_{i}\leq s\right\}\right]\right|\geq\frac{\sqrt{12}t}{4M_{1}\sqrt{rn}}\right)\leq\frac{1}{3t^{2}}.$
For the second piece, note the following:
$\left\lvert\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)^{2}-\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)^{2}\right\rvert\leq$ $2\left\lvert\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)\cdot\left\{\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)-\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)\right\}\right\rvert+$
|$\left\lvert\left\{\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)-\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)\right\}^{2}\right\rvert$
$\leq$ $2M_{1}\left\lvert\left\{\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)-\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)\right\}\right\rvert+$
|$\left\lvert\left\{\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)-\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)\right\}^{2}\right\rvert.$
Then for $c_{1}>0$ and $c_{2}>0$
$\Pr\left(\left\lvert\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)-\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)\right\rvert\geq\frac{c_{1}}{2M_{1}r}\right)\leq\frac{4M_{1}^{2}r^{2}\cdot M_{2}}{c_{1}^{2}n},$
and $\Pr\left(\left\lvert\left\{\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)-\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)\right\}^{2}\right\rvert\geq c_{2}\right)\leq\frac{M_{2}}{c_{2}n}.$
Once again, taking $c_{1}=\frac{2\sqrt{3}M_{1}rt\sqrt{M_{2}}}{\sqrt{n}}$ and $c_{2}=\frac{M_{2}t}{n}$:
$\Pr\left(\left\lvert\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)-\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)\right\rvert\geq\frac{t\sqrt{3M_{2}}}{\sqrt{n}}\right)\leq\frac{1}{3t^{2}},$
and $\Pr\left(\left\lvert\left\{\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)-\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)\right\}^{2}\right\rvert\geq\frac{M_{2}t}{n}\right)\leq\frac{1}{t}.$
Combining together, with probability $1-1/t-1/t^{2}$:
$\left\lvert\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i};X^{(k^{*})},s^{(k^{*})})\right]-\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X^{(k^{*})},s^{(k^{*})})\right]\right\rvert\leq\frac{K_{1}}{\sqrt{n}}+\frac{K_{2}}{n},$
where $K_{1}$ and $K_{2}$ are functions of $t$ and the constants $M_{1}$, $M_{2}$, and $M_{4}.$
B.3 Proof of Theorem 3.2
To prove Theorem 3.2, we will introduce a lemma that will be helpful in the following derivation. We define $n_{g}$ as the total number of units in a subgroup $g$, and $n_{g}(1)=\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})Z_{i}$ – i.e., $n_{g}(1)$ is equal to the number of units in subgroup $g$ who are treated. We similarly define $n_{g}(0)$ as the number of units in subgroup $g$ who are not treated.
[…]
  1. 1.
    Equiprobable treatment assignment patterns. For each subgroup $g\in\{1,\ldots,G\}$, all ($n_{g}\choose n_{g}(1)$ combination of ways to treat $n_{g}(1)$ units are equiprobable, given the subgroup size $n_{g}$.
  2. 2.
    Independent treatment assignment patterns: for all subgroups $g,g^{\prime}$, the treatment assignment process in group $g$ is independent of the treatment assignment process in group $g^{\prime}$, given $n_{g}(1)$ and $n_{g^{\prime}}(1)$.
We are considering desings that satisfy assignment symmetry. For example, experimental designs such as complete randomization, Bernoulli randomization, or blocking with independent treatment assignment across blocks, would satisfy assignment symmetry. However, cluster randomization would not satisfy assignment symmetry. Furthermore, we do not account for settings in which some units within a subgroup are more likely to receive treatment.
Lemma B.3 (Implications of Assignment Symmetry).
Under Assumption (B.1), we have assumed that conditional on the number of units within a subgroup $g\in\{1,\ldots,G\}$, the probability of receiving treatment for units in a subgroup is equiprobable and independent:
$\mathbb{E}\left\{Z_{i}\mid n_{g}(1)\right\}=\frac{n_{g}(1)}{n_{g}}\equiv\frac{1}{n_{g}}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})Z_{i},\text{ and }\mathbb{E}\left\{Z_{i}Z_{j}\mid n_{g}(1)=a\right\}=\frac{a(a-1)}{n_{g}(n_{g}-1)}$ (9)
We can then derive the distribution of $Z_{i}/n_{g}(1)$:
  • $\mathbb{E}\left[\frac{Z_{i}}{n_{g}(1)}\right]=\frac{1}{n_{g}}$
  • $\mathbb{E}\left[\frac{Z_{i}^{2}}{n_{g}(1)^{2}}\right]=\frac{1}{n_{g}}\mathbb{E}\left[\frac{1}{n_{g}(1)}\right]$
  • $\mathbb{E}\left[\frac{Z_{i}Z_{j}}{n_{g}(1)^{2}}\right]=\frac{1}{n_{g}(n_{g}-1)}\mathbb{E}\left[\frac{n_{g}(1)-1}{n_{g}(1)}\right]=\frac{1}{n_{g}(n_{g}-1)}\left(1-\mathbb{E}\left[\frac{1}{n_{g}(1)}\right]\right)$
  • $\mathbb{E}\left[\frac{Z_{i}(1-Z_{j})}{n_{g}(1)n_{g}(0)}\right]=\frac{1}{n_{g}\cdot(n_{g}-1)}\cdot\mathbb{E}\left[\frac{n_{g}-n_{g}(1)}{n_{g}(0)}\right]=\frac{1}{n_{g}\cdot(n_{g}-1)}$
The results of Lemma B.3 follow directly from first noting that conditional on the number of units in each subgroup $g$, randomization within a subgroup can be viewed as complete randomization, and then applying law of iterated expectations.
Part 2. Unbiasedness. To begin, we will show that $\hat{\tau}^{(g)}$ is unbiased for the within-sample subgroup ATE:
$\mathbb{E}$ $\left(\hat{\tau}^{(g)}\right)=\mathbb{E}\left\{\left.\frac{1}{\sum_{i=1}^{n}Z_{i}\hat{\mathcal{G}}_{g}(X_{i})}\sum_{i=1}^{n}Z_{i}Y_{i}\hat{\mathcal{G}}_{g}(X_{i})-\frac{1}{\sum_{i=1}^{n}(1-Z_{i})\hat{\mathcal{G}}_{g}(X_{i})}\sum_{i=1}^{n}(1-Z_{i})Y_{i}\hat{\mathcal{G}}_{g}(X_{i})~{}\right|\mathcal{D}\right\}$
$=\sum_{i=1}^{n}\mathbb{E}\left\{\left.\frac{Z_{i}}{\sum_{i=1}^{n}Z_{i}\hat{\mathcal{G}}_{g}(X_{i})}~{}\right|\mathcal{D}\right\}Y_{i}(1)\hat{\mathcal{G}}_{g}(X_{i})-\mathbb{E}\left\{\left.\frac{1-Z_{i}}{\sum_{i=1}^{n}(1-Z_{i})\hat{\mathcal{G}}_{g}(X_{i})}~{}\right|\mathcal{D}~{}\right\}Y_{i}(0)\hat{\mathcal{G}}_{g}(X_{i})$
Under Assumption (B.1) (Assignment Symmetry), we can apply the results from Lemma B.3:
$=\frac{1}{n_{g}}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})\left\{Y_{i}(1)-Y_{i}(0)\right\}$ (10)
[…]
Proof.
$\text{var}(\hat{\tau}^{(g)})$ $=\text{var}\left\{\frac{1}{\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})Z_{i}}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})Z_{i}Y_{i}-\frac{1}{\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})(1-Z_{i})}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})(1-Z_{i})Y_{i}\right\}$
$=\text{var}\left\{\frac{1}{n_{g}(1)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Z_{i}Y_{i}(1)-\frac{1}{n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}(1-Z_{i})Y_{i}(0)\right\}$
$\begin{split}&=\underbrace{\text{var}\left\{\frac{1}{n_{g}(1)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Z_{i}Y_{i}(1)\right\}}_{(1)}+\underbrace{\text{var}\left\{\frac{1}{n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}(1-Z_{i})Y_{i}(0)\right\}}_{(2)}\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-2\underbrace{\text{cov}\left\{\frac{1}{n_{g}(1)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Z_{i}Y_{i}(1),\frac{1}{n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}(1-Z_{i})Y_{i}(0)\right\}}_{(3)}\end{split}$ (11)
Beginning with term (1):
var $\left\{\frac{1}{n_{g}(1)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Z_{i}Y_{i}(1)\right\}$ (12)
$=\sum_{i,j:\hat{\mathcal{G}}_{g}(X)=1}\text{cov}\left(\frac{Z_{i}}{n_{g}(1)}Y_{i}(1),\frac{Z_{j}}{n_{g}(1)}Y_{j}(1)\right)$
$=\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left[\text{var}\left\{\frac{Z_{i}}{n_{g}(1)}Y_{i}(1)\right\}+\sum_{i\neq j}\text{cov}\left\{\frac{Z_{i}}{n_{g}(1)}Y_{i}(1),\frac{Z_{j}}{n_{g}(1)}Y_{j}(1)\right\}\right]$
$=\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left[\frac{1}{n_{g}}\mathbb{E}\left\{\frac{1}{n_{g}(1)}\right\}Y_{i}(1)^{2}-\frac{1}{n_{g}^{2}}Y_{i}(1)^{2}\right]+$
$\sum_{i\neq j}\frac{1}{n_{g}}\cdot\frac{1}{n_{g}-1}\cdot\left(1-\mathbb{E}\left\{\frac{1}{n_{g}(1)}\right\}\right)Y_{i}(1)Y_{j}(1)-\frac{1}{n_{g}^{2}}Y_{i}(1)Y_{j}(1)$
$=\frac{1}{n_{g}}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left(\mathbb{E}\left\{\frac{1}{n_{g}(1)}\right\}-\frac{1}{n_{g}}\right)Y_{i}(1)^{2}+$
$~{}~{}~{}~{}\frac{1}{n_{g}}\sum_{i\neq j:\hat{\mathcal{G}}_{g}(X)=1}\left(\frac{1}{n_{g}-1}\left(1-\mathbb{E}\left\{\frac{1}{n_{g}(1)}\right\}\right)-\frac{1}{n_{g}}\right)Y_{i}(1)Y_{j}(1)$
Let $\tilde{\beta}_{g}(z):=\mathbb{E}\{1/n_{g}(1)\}$:
$=\frac{1}{n_{g}}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left[\left\{\tilde{\beta}_{g}(1)-\frac{1}{n_{g}}\right\}Y_{i}(1)^{2}-\frac{1}{n_{g}-1}\sum_{i\neq j}\left\{\beta_{g}(1)-\frac{1}{n_{g}}\right\}Y_{i}(1)Y_{j}(1)\right]$
=$=\left\{\tilde{\beta}_{g}(1)-\frac{1}{n_{g}}\right\}\cdot\frac{1}{n_{g}}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left[Y_{i}(1)^{2}-\sum_{i\neq j}Y_{i}(1)Y_{j}(1)\right]$
=$=\left\{\tilde{\beta}_{g}(1)-\frac{1}{n_{g}}\right\}\text{var}_{g}\left\{Y_{i}(1)\right\}$ (13)
We can similarly show term (2) can be rewritten as:
$\text{var}\left\{\frac{1}{n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}(1-Z_{i})Y_{i}(0)\right\}=\left\{\beta_{g}(0)-\frac{1}{n_{g}}\right\}\text{var}_{g}\left\{Y_{i}(0)\right\}.$ (14)
Finally, to derive an expression for term (3), we begin by noting the following:
$\mathbb{E}$ [$\left[\left\{\frac{1}{n_{g}(1)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Z_{i}Y_{i}(1)\right\}\left\{\frac{1}{n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}(1-Z_{i})Y_{i}(0)\right\}\right]$
$=\mathbb{E}\left\{\frac{1}{n_{g}(1)\cdot n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Z_{i}(1-Z_{i})Y_{i}(1)Y_{i}(0)+\sum_{i\neq j}Z_{i}(1-Z_{j})Y_{i}(1)Y_{j}(0)\right\}$
Because $Z_{i}\cdot(1-Z_{i})=0$:
$=\mathbb{E}\left\{\frac{1}{n_{g}(1)\cdot n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\sum_{i\neq j}Z_{i}(1-Z_{j})Y_{i}(1)Y_{j}(0)\right\}$
$=\frac{1}{n_{g}\cdot(n_{g}-1)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\sum_{i\neq j}Y_{i}(1)Y_{j}(0)$
Then, we can re-write term (3) as:
cov $\left\{\frac{1}{n_{g}(1)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Z_{i}Y_{i}(1),\frac{1}{n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}(1-Z_{i})Y_{i}(0),\right\}$
$=\frac{1}{n_{g}}\frac{1}{n_{g}-1}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\sum_{i\neq j}Y_{i}(1)Y_{j}(0)-\left\{\frac{1}{n_{g}}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Y_{i}(1)\right\}\cdot\left\{\frac{1}{n_{g}}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Y_{i}(0)\right\}$
$=\frac{1}{n_{g}}\frac{1}{n_{g}-1}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\sum_{i\neq j}Y_{i}(1)Y_{j}(0)-\left[\frac{1}{n_{g}^{2}}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left\{Y_{i}(1)Y_{i}(0)+\sum_{i\neq j}Y_{i}(1)Y_{j}(0)\right\}\right]$
$=\frac{1}{n_{g}}\left(\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})}\frac{1}{n_{g}}Y_{i}(1)Y_{i}(0)+\left(\frac{1}{n_{g}-1}-\frac{1}{n_{g}}\right)\sum_{i\neq j}Y_{i}(1)Y_{j}(0)\right)$
$=\frac{1}{n_{g}}\left(\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})}\frac{1}{n_{g}}Y_{i}(1)Y_{i}(0)-\frac{1}{n_{g}(n_{g}-1)}\sum_{i\neq j}Y_{i}(1)Y_{j}(0)\right)$
$=-\frac{1}{n_{g}}\text{cov}_{g}(Y_{i}(1),Y_{i}(0))$ (15)
Substituting the expressions in Equation (13)-(15) into Equation (11):
$\text{var}(\hat{\tau}^{(g)})$ =$=\left\{\tilde{\beta}_{g}(1)-\frac{1}{n_{g}}\right\}\text{var}_{g}\left\{Y_{i}(1)\right\}+\left\{\tilde{\beta}_{g}(0)-\frac{1}{n_{g}}\right\}\text{var}_{g}\left\{Y_{i}(0)\right\}+\frac{2}{n_{g}}\text{cov}_{g}\left\{Y_{i}(1),Y_{i}(0)\right\}$
Defining $\beta_{g}(z):=n_{g}\tilde{\beta}_{g}(z)$ (i.e., the reciprocal of the proportion of units in group $g$ who are assigned $Z_{i}=z$):
$=\frac{\beta_{g}(1)}{n_{g}}\text{var}_{g}\left\{Y_{i}(1)\right\}+\frac{\beta_{g}(0)}{n_{g}}\text{var}_{g}\left\{Y_{i}(0)\right\}$
$~{}~{}~{}~{}~{}-\frac{1}{n_{g}}\left[\text{var}_{g}\left\{Y_{i}(1)\right\}+\text{var}_{g}\left\{Y_{i}(0)\right\}-2\text{cov}_{g}\left\{Y_{i}(1),Y_{i}(0)\right\}\right]$
$=\frac{\beta_{g}(1)}{n_{g}}\text{var}_{g}\left\{Y_{i}(1)\right\}+\frac{\beta_{g}(0)}{n_{g}}\text{var}_{g}\left\{Y_{i}(0)\right\}-\frac{1}{n_{g}}\text{var}_{g}(\tau_{i})$
B.4 Proof of Theorem 3.3
We formally impose the following regularity assumptions.
Assumption B.2 (Regularity Conditions).
  1. 1.
    As the sample size $n\to\infty$, the total number of subgroups $G$ remain fixed.
  2. 2.
    The proportion of units in each subgroup converges to a proportion $\pi_{g}^{*}\in(0,1)$ (i.e., $n_{g}/n\to\pi_{g}^{*}$, where $10<\pi_{g}^{*}<$ 1 for all $g\in\{1,\ldots,G\}$ and $\sum_{g=1}^{G}\pi_{g}^{*}=1$.
  3. 3.
    The proportion of treated units converges to $\pi^{*}_{z}$, where $\pi^{*}_{z}\in(0,1)$ (i.e., $\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}Z_{i}=\pi_{z}^{*}$).
  4. 4.
    Lindeberg Condition:
    $\lim_{n\to\infty}\frac{1}{n^{2}_{z}}\frac{max_{1\leq i\leq n}\hat{\mathcal{G}}_{g}(X_{i})Y_{i}^{2}(z)}{\text{var}(\hat{\tau}^{(g)})}=0.$
  5. 5.
    $\lim_{n\to\infty}\text{var}_{g}(Y_{i}(z))\leq c_{1},$ and $\text{cov}_{g}(Y_{i}(1),Y_{i}(0))\leq c_{2}$ where $c_{1},c_{2}<\infty$.
We would like to show that as $n\to\infty$, $\left(1-\frac{n_{z}}{n}\right)\cdot\frac{1}{n_{z}}\text{var}(\hat{\mathcal{G}}_{g}(X_{i}))\to 0$.
Claim: $\text{var}(\hat{\mathcal{G}}_{g}(X_{i}))=\frac{n}{n-1}\frac{1-\pi_{g}}{\pi_{g}}$.
Then,
$\left(1-\frac{n_{z}}{n}\right)\cdot\frac{1}{n_{z}}\text{var}(\hat{\mathcal{G}}_{g}(X_{i}))$ $=\frac{1}{n}\frac{1-n_{z}/n}{n_{z}/n}\cdot\frac{n}{n-1}\frac{1-\pi_{g}}{\pi_{g}}.$
As $n\to\infty$, $n_{z}/n\to\pi_{z}^{*}$ and $\pi_{g}\to\pi_{g}^{*}$, both of which will be bounded away from 0 and 1 (by Assumption (B.2)(b)-(c)):
$\lim_{n\to\infty}\left(1-\frac{n_{z}}{n}\right)\cdot\frac{1}{n_{z}}\text{var}(\hat{\mathcal{G}}_{g}(X_{i}))$ $=\lim_{n\to\infty}\frac{1}{n}\frac{1-p^{*}_{z}}{p^{*}_{z}}\cdot\frac{1-\pi^{*}_{g}}{\pi^{*}_{g}}=0$
[…]
$\frac{\hat{\tau}^{(g)}-\mathbb{E}\left(\hat{\tau}^{(g)}\right)}{\sqrt{\text{var}(\hat{\tau}^{(g)}})}\stackrel{{\scriptstyle d}}{{\rightarrow}}N(0,1)$
B.5 Proof of Corollary
Because each subgroup is a distinct partition, $\hat{\tau}^{(g)}$ and $\hat{\tau}_{g^{\prime}}$ will be asymptotically independent from one another. As a result, we can straightforwardly apply the Cramer-Wold device and
$\boldsymbol{\hat{\tau}}-\boldsymbol{\tau}\stackrel{{\scriptstyle d}}{{\rightarrow}}N(0,\boldsymbol{\Sigma}_{G})$
Thus:
$\boldsymbol{\hat{\tau}}^{\top}\boldsymbol{\Sigma}^{-1}_{G}\boldsymbol{\hat{\tau}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\chi^{2}_{G}$
B.6 Proof of Example 3.1
Assumption B.3 (Regularity Conditions).
  1. (a)
    Smoothness of $\tau(X)$: $\tau(X)$ is continuous, and $\tau^{\prime}(X)$, $\tau^{\prime\prime}(X)$ exist and are uniformly bounded in a neighborhood around $s$, and $\tau^{\prime}(s)\neq 0$.
  2. (b)
    Smoothness condition on density of $X$ and $v$: $dF^{\prime}_{X}$ exists and is uniformly bounded in a neighborhood around $s$. Furthermore, $dF_{X}(s)\neq 0$.
  3. (c)
    Moment condition: $\mathbb{E}(v_{i})=0$, $\mathbb{E}(v_{i}^{2})=\sigma^{2}$.
  4. (d)
    Tail condition: $dF_{\tau}=o(|\tau|^{-4+\delta})$ as $|\tau|\to\infty$ for some $\delta>0$.
  5. (e)
    Signal condition: $\text{var}(\tau(X_{i}))>0$.
[…]
Lemma B.4 (Asymptotic Variance of Splits).
Under Assumption (B.3) and squared loss, the asymptotic variance of the estimated splits is:
[…]
where $V=-dF_{X}(s)f^{\prime}(s)\neq 0$.
Proof.
[…]
$n^{1/3}(\hat{s}_{n}-s)\stackrel{{\scriptstyle d}}{{\rightarrow}}W_{\sigma^{2},s}:=\operatorname*{arg\,max}_{t}\left[Q(t)\cdot\text{sign}\left(\mathbb{E}(\tau_{i}\mid X_{i}<s)-\mathbb{E}(\tau_{i}\mid X_{i}\geq s)\right)\right],$
and $Q(t)$ is a scaled, two-sided Brownian motion, originating from zero, with a quadratic drift:
$Q(t)=dF_{X}(s)\sigma^{2}B(t)-\frac{1}{2}Vt^{2},$
[…]
To begin, define the random variable $Z$ as:
$Z_{\gamma}:=\operatorname*{arg\,max}_{t}B(t)-\gamma t^{2},$
[…]
Furthermore, in settings when the Brownian motion is scaled (i.e., $B(at)$), we can exploit the fact that for all $a>0$, $B(at)=a^{1/2}B(t)$, and:
$Z_{\gamma}=^{d}a\operatorname*{arg\,max}(a^{1/2}B(t)-a^{2}\gamma t^{2})=aZ_{a^{3/2}\gamma}.$
We then solve for the variance of $W_{\sigma^{2},s}$ by substituting $a^{1/2}=dF_{X}(s)\sigma^{2}$, and $\gamma=\frac{1}{2}V/(dF_{X}(s)^{2}\sigma^{4})^{2}$. Thus, the asymptotic variance of a split $\hat{s}_{n}$ can be written as follows:
[…]
B.7 Proof of Lemmas
B.7.1 Proof of Lemma B.1
Proof.
We can re-write the difference in the within-sample loss and the population loss as:
$\left\lvert\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i},X;\hat{s}_{n})\right]-\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X,s)\right]\right\rvert$
$=\left\lvert\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i},X;\hat{s}_{n})\right]-\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i};X,s)\right]+\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i};X,s)\right]-\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X,s)\right]\right\rvert$
$\leq\underbrace{\left\lvert\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i},X;\hat{s}_{n})\right]-\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i};X,s)\right]\right\rvert}_{(*)}+\left\lvert\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i};X,s)\right]-\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X,s)\right]\right\rvert$
We will construct an upper bound on the first term ()(*)( ∗ ):
$\left|\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i};X,\hat{s}_{n})\right]-\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i};X,s)\right]\right|$
$=\left|\mathbb{E}_{n}\left[(\hat{\tau}^{d}_{i}-\tau^{d}(X_{i};\hat{s}_{n}))^{2}-(\hat{\tau}^{d}_{i}-\tau^{d}(X_{i};s))^{2}\right]\right|$
$=\left|\mathbb{E}_{n}\left[\tau^{d}(X_{i};\hat{s}_{n})^{2}-\tau^{d}(X_{i};s)^{2}-2\hat{\tau}^{d}_{i}\left(\tau^{d}(X_{i};\hat{s}_{n})-\tau^{d}(X_{i};s)\right)\right]\right|$
$=\left|\mathbb{E}_{n}\left[\left(\tau^{d}(X_{i};\hat{s}_{n})-\tau^{d}(X_{i};s)\right)\left(\tau^{d}(X_{i};\hat{s}_{n})+\tau^{d}(X_{i};s)-2\hat{\tau}^{d}_{i}\right)\right]\right|$
$\leq\mathbb{E}_{n}\left[\left|\left(\tau^{d}(X_{i};\hat{s}_{n})-\tau^{d}(X_{i};s)\right)\right|\left|\left(\tau^{d}(X_{i};\hat{s}_{n})+\tau^{d}(X_{i};s)-2\hat{\tau}^{d}_{i}\right)\right|\right]$
$=\mathbb{E}_{n}\left[\left|\left(\mathbb{E}(\hat{\tau}^{d}_{i}|X_{i}<\hat{s}_{n})-\mathbb{E}(\hat{\tau}^{d}_{i}|\ X_{i}<s)\right)\right|\left|\left(\tau^{d}(X_{i};\hat{s}_{n})+\tau^{d}(X_{i};s)-2\hat{\tau}^{d}_{i}\right)\right|\right]$
$=\mathbb{E}_{n}\left[\left|\hat{s}_{n}-s\right|\times\max_{c\in[\hat{s}_{n},s]}\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}<c)\left|\tau^{d}(X_{i};\hat{s}_{n})+\tau^{d}(X_{i};s)-2\hat{\tau}^{d}_{i}\right|\right]$
$=\left|\hat{s}_{n}-s\right|\mathbb{E}_{n}\left[\max_{c\in[\hat{s}_{n},s]}\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}<c)\left|\left(\tau^{d}(X_{i};\hat{s}_{n})-\hat{\tau}^{d}_{i}\right)+\left(\tau^{d}(X_{i};s)-\hat{\tau}^{d}_{i}\right)\right|\right]$
$\leq\left|\hat{s}_{n}-s\right|\times\max_{c\in[\hat{s}_{n},s]}\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}<c)\times\left(\mathbb{E}_{n}\left|\tau^{d}(X_{i};\hat{s}_{n})-\hat{\tau}^{d}_{i}\right|+\mathbb{E}_{n}\left|\tau^{d}(X_{i};s)-\hat{\tau}^{d}_{i}\right|\right)$
$\leq 2\left|\hat{s}_{n}-s\right|\times\max_{c\in[\hat{s}_{n},s]}\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}<c)\times\mathbb{E}_{n}\left|\tau^{d}(X_{i};\hat{s}_{n})-\hat{\tau}^{d}_{i}\right|$
$=C_{\tau}\times\left\lvert\hat{s}_{n}-s\right\rvert$
Then:
$\left\lvert\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i},X;\hat{s}_{n})\right]-\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X,s)\right]\right\rvert\leq C_{\tau}\times\left\lvert\hat{s}_{n}-s\right\rvert+\bigg{\lvert}\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i};X,s)\right]-\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X,s)\right]\bigg{\rvert},$
which concludes the proof.
B.8 Proof of Lemma B.2
Proof.
We have that
$\Pr$ $\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}_{n}^{(b)})\right\}<\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}\right]$
$=\Pr\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}^{(b)}_{n})\right\}-\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}<0\right]$
$=\Pr\Big{[}\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}+$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}-\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},s^{(a)})\right\}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}+\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},\hat{s}^{(a)}_{n})\right\}<0\Big{]}$
By construction, $\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}>\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}$, as $X^{(a)}$ is a relevant covariate. Therefore, $\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},s^{(a)})\right\}=\delta_{ba}>0$:
$=\Pr\left(\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}-\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right]+\delta_{ba}<0\right)$
$=\Pr\left(\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right]\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.-\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}\right]>\delta_{ba}\right)$
$\leq\Pr\left(\left|\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right]\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\left.\left.\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}_{i};X^{(b)},s^{(b)})\right\}\right]\right|>\delta_{ba}\right)$
Then, applying Markov’s Inequality:
$\leq\frac{1}{\delta_{ba}}\mathbb{E}\left(\left|\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right]\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.-\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},s^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}\right]\right|\right)$
$\leq\frac{1}{\delta_{ba}}\left\{\mathbb{E}\left[\left|\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right|\right]\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.+\mathbb{E}\left[\left|\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}\right|\right]\right\}$
Without loss of generality, assume $\mathbb{E}\left[\left|\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right|\right]$ is greater than or equal to $\mathbb{E}\left[\left|\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}^{(b)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(b)},s^{(b)})\right\}\right|\right]$. Then:
$\leq\frac{2}{\delta_{ba}}\mathbb{E}\left[\left|\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}^{(a)}_{n})\right\}-\mathbb{E}\left\{\ell(\hat{\tau}^{d}_{i};X^{(a)},s^{(a)})\right\}\right|\right]$
Appendix C Method Implementation Details
[…]
Causal Distillation Trees (CDT).
[…]
To perform honest estimation of the subgroup ATEs, we set $\pi_{train}=0.70$ for splitting the data into a training and hold-out estimation set. To estimate the individual-level treatment effects $\hat{\tau}(X_{i})$, we used the out-of-bag sample estimates (without repeated cross-fitting) in the Distilled Causal Forest and Distilled BCF, and we used the repeated cross-fitting procedure with $R=50$ repeats in the Distilled Rboost. We also examined other choices of $R$ in Figure A8. From these simulations, we found that $R=50$ provided a good balance between the subgroup estimation accuracy and computational burden. However, more generally, the subgroup estimation results from Distilled Rboost were robust as long as $R$ was sufficiently large enough.
Virtual Twins.
[…]
Causal Trees.
[…]
Linear and Lasso Regression.
[…]
Appendix D Simulations
D.1 Details on Simulation Evaluation Metrics
To evaluate the effectiveness of each subgroup estimation method, we consider the accuracy of subgroup identification from three perspectives:
  1. 1.
    Selected subgroup features: Whether the features used to define the estimated subgroups match the features used to define the true subgroups (i.e., $X^{(1)}$ and $X^{(2)}$), as measured by the accuracy of the selected subgroup features, we calculate the number of true positives, false positives, and $F_{1}$ score. [^7] The $F_{1}$ score summarizes the number of true positives (TP), false positives (FP), and false negatives (FN) into a single quantity (i.e., $F_{1}=\frac{2\cdot TP}{2\cdot TP+FP+FN}$)
  2. 2.
    Estimated subgroup thresholds: Whether the estimated subgroup thresholds are close to the true subgroup thresholds, as measured via the root mean squared error (RMSE).
  3. 3.
    Estimated Subgroup ATEs: Whether the estimated subgroup average treatment effects $\hat{\tau}^{(g)}$ are close to the true subgroup average treatment effects (i.e., $\tau^{(g)}:=\mathbb{E}\left[\tau_{i}\mid\mathcal{G}_{g}(X_{i})\right]$), as measured via the RMSE.
D.2 Additional Simulation Settings
We next present additional simulation results to complement the main results, shown in Section 5. These results include additional subgroup data-generating processes, evaluation metrics, and alternative modeling choices in the CDT framework (e.g., different number of repeated cross-fits $R$, student model choices, and pruning choices).
CATE-only Outcome Model.
In particular, to complement Figure 3 which examined the subgroup estimation performance under an outcome model including linear covariate effects (i.e., $Y_{i}=Z_{i}\cdot\tau_{i}+X_{i}^{(3)}+X_{i}^{(4)}+\nu_{i}$), we also examined the subgroup estimation performance under the following outcome model, which depends only on the CATE:
$Y_{i}=Z_{i}\cdot\tau_{i}+\nu_{i}.$
Besides this modification in the outcome model, all other simulation settings were kept the same. The results are shown in Figure A3. As seen under the previous outcome model, the distilled methods (Distilled Causal Forest and Distilled Rboost) generally outperformed existing methods in estimating the subgroup structure and the subgroup ATEs. Moreover, the performance of the linear models for estimating the subgroup ATEs significantly deteriorated under this CATE-only outcome model. This highlights the sensitivity of the linear model approaches to the form of the outcome model. In comparison, CDT tends to work well in both the setting considered here and in Section 5.
Figure A3: Performance of subgroup estimation methods for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes in the CATE-only outcome model scenario (rows). CDT with various teacher models (i.e., Distilled Causal Forest and Distilled Rboost) frequently yields the highest $F_{1}$ and number of true positives alongside the lowest number of false positives, threshold RMSEs, and subgroup ATE RMSE, demonstrating its effectiveness for accurate subgroup estimation. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
Threshold Distributions.
When examining the distribution of the estimated subgroup thresholds from the tree-based methods in Figure A4, we uncover another benefit of distillation — the distribution of the CDT-estimated subgroup thresholds using Rboost or causal forest is much tighter (i.e,. has smaller variance) than for other tree-based subgroup detection methods without distillation (i.e., causal trees and virtual twins). This again reinforces our theoretical understanding of CDT, where we have seen in Example 3.1 that the first-stage learner in CDT acts as a de-noising step, leading to more stable splits (i.e., thresholds) in CDT compared to non-distilled tree-based methods.
Note that while the thresholds distribution from Distilled BCF is more variable than the non-distilled methods in weak-to-moderate treatment effect heterogeneity regimes, Distilled BCF would have been easily ruled out in these settings according to our teacher model selection procedure using Jaccard SSI (see Appendix E).
Figure A4: Distribution of estimated thresholds for each true subgroup feature (i.e., $X^{(1)}$ and $X^{(2)}$) using different subgroup estimation methods (color) under various treatment effect heterogeneity strengths (columns) and subgroup data-generating processes (subplots). The range of the threshold distributions from Distilled Causal Forest and Distilled Rboost is often smaller than that from causal tree, especially as the treatment effect heterogeneity strength increases. Results are shown for 100 simulation replicates.
Pruned versus Unpruned.
Recall thus far that we have been implementing CDT and causal trees with the standard post-pruning procedure using cross-validation to tune the complexity parameter $\alpha$. To investigate the impact of this post-pruning step, we compare the subgroup estimation performance of CDT and causal trees with and without post-pruning in Figure A5. While the performance of causal trees declines drastically without pruning, the performance of CDT using causal forest or Rboost is similarly strong with and without pruning, demonstrating their robustness with respect to this pruning choice. For Distilled BCF, pruning improves the subgroup estimation performance, particularly in the weak-to-moderate treatment effect heterogeneity regimes.
Figure A5: Performance of pruned and unpruned versions of CDT and causal trees for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes (rows). While causal tree without pruning performs substantially worse than pruned causal tree, CDT with various teacher models (e.g., Distilled Causal Forest and Distilled Rboost) performs similarly well with or without pruning. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
Evaluating using Oracle Tree Depth.
[…]
Figure A6: Performance of subgroup estimation methods with oracle pruning for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes (rows). CDT with various teacher models remains the most robust and accurate subgroup estimation method, as seen by its high $F_{1}$ and number of true positives alongside low number of false positives, threshold RMSEs, and subgroup ATE RMSE. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
Selection Frequency of Subgroup Features.
In Figure A7, we show the number of simulation replicates (out of 100) that each variable was selected as a subgroup feature. Notably, when the outcome model includes linear covariate effects involving the features $X^{(3)}$ and $X^{(4)}$, causal tree and virtual twins frequently split on these irrelevant features, $X^{(3)}$ and $X^{(4)}$, whereas the distilled methods avoid this pitfall.
Figure A7: Number of simulation replicates out of 100, for which each variable was selected (at least once) in the estimated subgroups. Results are shown for different variables (x-axis), subgroup estimation methods (y-axis), treatment effect heterogeneity strengths (rows), and subgroup data-generating processes (subplots). Large values for variables $X^{(1)}$ and $X^{(2)}$ and small values for all other variables indicate a better-performing method.
Choice of Repeated Cross-Fits $R$.
In the CDT algorithm, we have been using $R=50$ repeated cross-fits to estimate the individual-level treatment effects. To investigate the sensitivity of this choice, we show in Figure A8 the subgroup estimation performance of Distilled Rboost for varying choices of $R$, ranging between $1$ and $100$. In general, the subgroup estimation performance improves as $R$ increases and tends to be relatively stable for $R>10$.
Figure A8: Performance of Distilled Rboost with varying number of repeated crossfits $R$ (color) for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes (rows). The subgroup estimation performance improves slightly as $R$ increases and tends to be relatively stable for $R>10$. Results are averaged across 100 simulation replicates.
Rulefit Student Model.
[…]
  1. 1.
    Rulefit (rules only, max depth = 2): type = "rules" and maxdepth = 2
  2. 2.
    Rulefit (rules only, max depth = 3): type = "rules" and maxdepth = 3
  3. 3.
    Rulefit (linear + rules, max depth = 2): type = "both" and maxdepth = 2
  4. 4.
    Rulefit (linear + rules, max depth = 3): type = "both" and maxdepth = 3 (i.e., the default settings in pre::pre())
Although the rulefit student model sometimes results in more accurate estimation of the subgroup ATE, this is at the cost of a more complex model, illustrated by the substantially higher number of false positive features in rulefit compared to CART. CART thus appears to be a simpler and more interpretable student model choice for CDT.
Figure A9: Performance of CDT with CART versus Rulefit student models for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes (rows). The Rulefit student model tends to estimate unnecessarily complex subgroups, compared to CART. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
Appendix E Additional Teacher Model Selection Simulation Results
In this section, we provide the Jaccard SSI simulation results (Figures A10-A15), used to select the teacher model, for each of the subgroup DGPs studied in Section 5 and Appendix D. As discussed in Section 5, higher Jaccard SSI generally corresponds to more accurate subgroup estimation regardless of the choice of subgroup DGP.
In practice, we reiterate that substantive researchers should leverage their domain knowledge when choosing the tree depth(s) $d$ to consider in the teacher model selection procedure. In cases where such prior knowledge is limited, we recommend that researchers view the Jaccard SSI results alongside the feature stability distributions and other diagnostic tools (see Appendix A.8) to gain a more holistic perspective. This holistic view of the distilled method and its stability can often shed light on an appropriate choice of tree depth. For example, the feature stability distributions in Figures A10-A15 clearly illuminate that the trees grown to depths 3 and 4 are substantially more unstable than the depth-2 trees. That is, the distribution of features selected at depths 3 and 4 are far more heterogeneous than that at depth 2. As with the Jaccard SSI, a more stable or homogeneous feature distribution is generally a positive sign and may serve as heuristic to help choose the tree depth.
Figure A10: Under the ‘AND’ subgroup data-generating process (CATE-only), we examine (A) the subgroup estimation accuracy alongside (B) the Jaccard SSI for a range of tree depths and (C) the distribution of features selected at each tree depth across the 100 bootstraps. Results are shown for different subgroup estimation methods (colors) and treatment effect heterogeneity strengths (rows). Choosing the teacher model in CDT which leads to the highest Jaccard SSI generally corresponds to more accurate subgroup estimation. Moreover, the amount of heterogeneity in the distribution of selected subgroup features can help inform our degree of trust and guide selection for choosing the relevant tree depth(s).
Figure A11: Under the ‘Additive’ subgroup data-generating process (CATE-only), we examine (A) the subgroup estimation accuracy alongside (B) the Jaccard SSI for a range of tree depths and (C) the distribution of features selected at each tree depth across the 100 bootstraps. Results are shown for different subgroup estimation methods (colors) and treatment effect heterogeneity strengths (rows). Choosing the teacher model in CDT which leads to the highest Jaccard SSI generally corresponds to more accurate subgroup estimation. Moreover, the amount of heterogeneity in the distribution of selected subgroup features can help inform our degree of trust and guide selection for choosing the relevant tree depth(s).
Figure A12: Under the ‘OR’ subgroup data-generating process (CATE-only), we examine (A) the subgroup estimation accuracy alongside (B) the Jaccard SSI for a range of tree depths and (C) the distribution of features selected at each tree depth across the 100 bootstraps. Results are shown for different subgroup estimation methods (colors) and treatment effect heterogeneity strengths (rows). Choosing the teacher model in CDT which leads to the highest Jaccard SSI generally corresponds to more accurate subgroup estimation. Moreover, the amount of heterogeneity in the distribution of selected subgroup features can help inform our degree of trust and guide selection for choosing the relevant tree depth(s).
Figure A13: Under the ‘AND’ subgroup data-generating process with linear covariate effects, we examine (A) the subgroup estimation accuracy alongside (B) the Jaccard SSI for a range of tree depths and (C) the distribution of features selected at each tree depth across the 100 bootstraps. Results are shown for different subgroup estimation methods (colors) and treatment effect heterogeneity strengths (rows). Choosing the teacher model in CDT which leads to the highest Jaccard SSI generally corresponds to more accurate subgroup estimation. Moreover, the amount of heterogeneity in the distribution of selected subgroup features can help inform our degree of trust and guide selection for choosing the relevant tree depth(s).
Figure A14: Under the ‘Additive’ subgroup data-generating process with linear covariate effects, we examine (A) the subgroup estimation accuracy alongside (B) the Jaccard SSI for a range of tree depths and (C) the distribution of features selected at each tree depth across the 100 bootstraps. Results are shown for different subgroup estimation methods (colors) and treatment effect heterogeneity strengths (rows). Choosing the teacher model in CDT which leads to the highest Jaccard SSI generally corresponds to more accurate subgroup estimation. Moreover, the amount of heterogeneity in the distribution of selected subgroup features can help inform our degree of trust and guide selection for choosing the relevant tree depth(s).
Figure A15: Under the ‘OR’ subgroup data-generating process with linear covariate effects, we examine (A) the subgroup estimation accuracy alongside (B) the Jaccard SSI for a range of tree depths and (C) the distribution of features selected at each tree depth across the 100 bootstraps. Results are shown for different subgroup estimation methods (colors) and treatment effect heterogeneity strengths (rows). Choosing the teacher model in CDT which leads to the highest Jaccard SSI generally corresponds to more accurate subgroup estimation. Moreover, the amount of heterogeneity in the distribution of selected subgroup features can help inform our degree of trust and guide selection for choosing the relevant tree depth(s).
Appendix F Extended Details of Case Study
F.1 Applying Jaccard SSI
While we presented the final subgroups and their estimated effects under Distilled Rboost in the main text, this is only after comparing our proposed Jaccard SSI on across two different teacher models. These were Causal Forest and Rboost. Figure A16 indicates that Distilled Rboost is consistently more stable across different tree depths.
Figure A16: In our case study, we examine the the Jaccard SSI for a range of tree depths. Results are shown for different subgroup estimation methods (colors). Choosing the teacher model in CDT which leads to the highest Jaccard SSI generally corresponds to more accurate subgroup estimation.
F.2 Analyzing the subgroup treatment effects
[…]
Subgroup Results
[…]
Table 1: We provide the average treatment effect across subgroups estimated by CDT using Rboost as the teacher model. CDT is trained on 50% of the study and treatment effects are estimated on the remaining 50% for honest estimation. Bootstrapped standard errors are reported in parentheses.
F.3 Stability analysis
To consider the stability of the different methods, we perform a bootstrap analysis. Across 100 different bootstrap samples, we apply CDT, virtual twins, as well as tree-based methods and linear methods, to compare the different subgroups obtained under each technique. For each bootstrap instance, we compute the proportion of times features were used to construct subgroups. We summarize the results across these runs in Figure A17.
From the bootstrap analysis, we find that the distillation methods consistently split on the same set of features, regardless of the data perturbations. Furthermore, CDT resulted in fewer subgroups with lower variability, and the estimated subgroups were overwhelmingly characterized by features with direct clinical relevance. Consistent with the one-shot analysis, we see that the three main features chosen by the CDT methods were CD8 cell count at baseline, participants weight at the start of the study, and the number of days of pre-175 antiretroviral therapy.
Figure A17: For our case study, we summarize the proportion of times across the 100 bootstrap instances features are split upon to form rules and compose subgroups across methods (colors). Proportions closer to one indicate more stability and robustness to small data perturbations.
Causal tree also similarly constructs subgroups using CD8 cell count, participants’ starting weight, and the number of days of pre-175 antiretroviral therapy. However, it additionally splits on other covariates, depending on the bootstrap iteration, with less consistency. Similarly, virtual twins consistently constructs subgroups on those same three features, but also constructs substantially more subgroups using other features, resulting in more subgroups that are challenging to interpret.
[…]